### v1

# Constant-Size Certificates for Leader Election in Chordal Graphs and Related Classes

Abstract
In distributed computing a certification scheme consists of a set of states and conditions over those states that enable each node of a graph to efficiently verify the correctness of a solution to a given problem. This work focuses on two fundamental problems: leader election and spanning tree construction. For each problem, we present a constant-size (per edge), local certification scheme, where the conditions available to each node can only refer to the graph induced by its one-hop neighborhood. In particular, we provide certification schemes for leader election in chordal and $K_{4}$-free dismantlable graphs and for spanning tree construction in dismantlable graphs, assuming a root is given. For chordal graphs, our leader election certification scheme additionally ensures an acyclic orientation, a property that is not generally verifiable using constant-size certificates in arbitrary graphs. To the best of our knowledge, these are the first local certification results tailored to these graph classes, potentially highlighting structural properties useful for verifying additional problems. Finally, we propose an algorithm that automatically transforms any certification scheme into a silent self-stabilizing algorithm (i.e., an algorithm that automatically recovers from faults) by adding only one extra state to the set of states of the certification scheme, assuming a Gouda fair scheduler. This transformation may be of independent interest.
keywords:
Local Certification, Distributed Verification, Proof Labelling Schemes, Chordal Graphs, Dismantlable Graphs, Leader Election, Distributed Computing
1 Introduction
Local certification [feuilloley2019introduction], also known as proof labelling schemes [korman2005proof], locally checkable proofs [goos2016locally] or distributed verification, verifies global properties through local conditions. Each node individually assesses these conditions and either accepts (i.e., all conditions are satisfied) or rejects (i.e., at least one condition is violated). An error is detected if any node rejects. For example, to verify that a graph is a cycle and not a path, each node accepts if its degree is two and rejects if its degree is one. However, the inverse problem of accepting a path and rejecting a cycle is not possible with certificates of constant size, even if the nodes are not identical [feuilloley2019introduction].
The ability to continue functioning and recovering from faults is essential for distributed systems. Dijkstra [dijkstra1974stabilising] introduced the concept of self-stabilization as a generic way to model various faults and guarantee the correct behavior of an algorithm regardless of transient faults. An algorithm is self-stabilizing if it eventually converges to a correct output from any arbitrary initial configuration (see also [dolev2000book]). A self-stabilizing algorithm is called silent if when it reaches a correct configuration, it remains in that configuration unless a fault occurs. On the other hand, a self-stabilizing algorithm that is not silent can transition between correct configurations as the algorithm perpetually continues computing. The design of a self-stabilizing algorithm typically consists of three components: (a) Detect when the global configuration is faulty (b) Reach a non-faulty configuration (c) Solve the original task. The first aspect naturally connects self-stabilization with local certification, where nodes maintain local certificates that allow verification of the global configuration via local checks. For instance, in the self-stabilizing spanning tree algorithm of [chen1991spanning], each node stores a pointer to its parent (as a way to solve the task) and its distance from the root (as a way to verify the solution). Without the latter, cycles cannot be locally detected.
As a form of fault-tolerance, verification is not the main task to be solved within a system and as a result it should not consume too many resources. Hence, in this work, we only consider certificates of constant size per edge. A significant challenge in this setting is that in general problems such as the detection of cycles mentioned previously require more than constant space. However, the existence of models such as Population Protocols [aspnes2009population] and Stone Age Computing [emek2013stoneage] which only use constant space demonstrates that such constraints do not preclude the existence of solutions to non-trivial problems.
The main problem that we consider in the following sections is Leader Election [LeLann77], which is an important primitive in distributed systems as it allows the coordination of processes in the system. The goal in this problem is for a unique node to enter and remain in the elected state indefinitely. A usual assumption within this context, that we do not make in this work, is that nodes have distinct identifiers. In the anonymous setting, the problem cannot be solved without additional assumptions due to symmetries [angluin1980local]. One way to overcome this impossibility is to characterize graph classes where the problem can be solved (e.g., as in [yamashita1996computing]) or to employ randomization (e.g., [itai1990symmetry]). In this work, we use the former method and provide certificates for leader election for the case of chordal and $K_{4}$-free dismantlable graphs.
Both chordal graphs and dismantlable graphs can be characterized by an ordering of their vertices. Chordal graphs admit an ordering called a simplicial elimination scheme, in which each vertex is adjacent to a clique of nodes that appear later in the ordering. Similarly, dismantlable graphs admit a dismantling order, where each vertex $v$ has a neighbor that is adjacent to all other neighbors of $v$ appearing later in the ordering. Dismantlability, in particular, is a property shared by many important classes in metric graph theory, such as bridged [FaJa, SoCh] and Helly graphs [BP-absolute, BaPr]. In this work, we design certificates that take advantage of these orderings to define and locally verify an orientation of the edges incident to each node, ensuring that the graph contains a unique sink node, which we define to be the leader.
1.1 Related Work
In the LOCAL model introduced by Linial [linial1987local, linial1992journal], it is assumed that nodes exchange messages with neighboring nodes until a task is solved. The complexity of an algorithm within this model thus depends on the size of the messages and on the number of synchronous communication rounds. An alternative definition of the model that is more appropriate for our setting, is to assume that every node can see the state of each node in its neighborhood up to some distance $d$ and based on this partial knowledge of the graph, choose an output. The goal in this case is to minimize $d$ and the size of states for some given task. What can be computed within this setting has received significant attention (e.g., [vishkin1986determinisitic, goldberg1987parallel, luby1986simple, alon1986fast]). The concept of locally verifiable problems was formalized by Naor and Stockmeyer in their seminal paper [naor1993lcl, naor1995journal] in which they defined the family of Locally Checkable Labeling (LCL) problems. A problem is in LCL if given the view, consisting of constant-sized input and output labels, at some constant distance $d$ for each node, there exists an algorithm that can decide whether a computed solution is correct. Although LCL problems were continually explored since their introduction, no groundbreaking results emerged until 2016 when Brandt et al. [brandt2016lower] led to a series of papers (e.g., [balliu2018classes, balliu2018locally, balliu2020randomness]) that gradually provided a complete picture on the classification of the round complexity of all LCL problems in four broad classes [suomela2020landscape].
Recently, Bousquet et al. [feuilloley2025certification] proposed a different direction of classification based on space instead of round complexity. Several results are given in [feuilloley2025certification], characterizing both “gaps” (i.e., pairs of functions $f(n),g(n)$, with $f(n)<g(n)$ where no property has optimal certificate size in $(f(n),g(n))$ for a given class of graphs) as well as proving the existence of properties with optimal certificate sizes for various classes of graphs including paths, cycles and caterpillars. In this work we consider chordal and ($K_{4}$-free) dismantlable graphs which are not explicitly studied in [feuilloley2025certification]. Our approach is complementary to that of [feuilloley2025certification] in the sense that we only consider specific problems but for larger classes of graphs.
The most similar paper to our work is [chalopin2024stabilising]. Therein, the authors propose a certificate and a set of conditions with constant memory per node, ensuring a unique sink for a simply connected subset of the regular triangular grid, where simply connected means that the configuration does not contain holes. Observe that simply connected parts of triangular grids are $K_{4}$-free dismantlable graphs, hence this work (in particular Section˜4) generalizes the result of [chalopin2024stabilising] and extends this idea to additional classes of graphs. Additionally, our definition of view is similar to the concept of binoculars introduced in [chalopin2015binoculars]. Binoculars provide each node $v$ with all information on the graph induced by the neighborhood of $v$ at distance one.
1.2 Our Contributions
We define constant-size certification schemes for verifying the existence of a unique sink in chordal graphs and $K_{4}$-free dismantlable graphs, under the assumption that each node can see the subgraph induced by its one-hop neighbors and the certificates of its neighbors. This notion of view can be seen as an intermediate assumption, requiring more information than standard radius one visibility but without accessing nodes at distance two. For chordal graphs, the certification scheme also ensures a valid acyclic orientation, which is particularly important as acyclicity cannot, in general, be verified with constant-size certificates. In the case of dismantlable graphs, we also provide a certification scheme to ensure the existence of a spanning tree, assuming there is a distinct node marked as the root. The certification scheme verifying the existence of a unique sink in $K_{4}$-free dismantlable graphs can be used to verify the existence of such a root. All of our certificates require only constant size per edge.
In the last section, we show that given a constant-size certificate, a silent self-stabilizing algorithm can always be constructed with constant memory per process, assuming a Gouda fair scheduler. Let $k$ be the number of bits needed to verify a task $T$ using a certificate. Our algorithm transforms the certificate to a self stabilizing algorithm using $k+1$ bits. This is less than prior general transformations, such as the one in [BlinFP14], which requires $O(k+\log n)$ bits, where $n$ is the number of computational entities in the system, but works even under an unfair scheduler. As a result, Algorithm 1 can be used in systems with constant memory where [BlinFP14] cannot be employed.
Our results demonstrate that self-stabilization with strong global guarantees (i.e., unique sink, acyclic orientation) is possible with limited local memory in natural graph classes. That is, we show that combining certificate-based verification and Gouda fairness provides a way for designing fault-tolerant systems even when only constant memory is available.
2 Model and Preliminaries
2.1 Model
Let $G=(V,E,\delta)$ be a graph with vertex set $V$, edge set $E$ and a port labeling function $\delta$. For any vertex $v\in V$, let $N(v)$ denote its neighborhood and let $N[v]=N(v)\cup\{v\}$ denote its closed neighborhood. The 1-ball at $v$, denoted by $B_{1}(v)$, is the graph induced by $N[v]$. When $G$ is oriented, we write $N^{+}(v)$ (resp. $N^{-}(v)$) for the outgoing (resp. incoming) neighborhood of $v$ in $G$ and refer to nodes in $N^{+}(v)$ as outneighbors of $v$. We define the port labeling function to be $\delta=\{\delta_{v}\}_{v\in V}$, where $\delta_{v}:N(v)\rightarrow\{1,2,\ldots,deg(v)\}$ assigns a unique port number to each edge incident to $v$ for every $v\in V$ (i.e., $\delta_{v}$ is a bijection). We write $K_{4}$ to denote a complete graph with four vertices. A graph is $K_{4}$-free if it contains no subgraph that is isomorphic to a $K_{4}$. We write directed triangle to refer to a cycle of size three.
Let $G=(V,E,\delta)$ be a graph and $S$ be an arbitrary set of states. For any $G$ and any mapping $x_{S}:V\rightarrow S$, we say that $(G,x_{S})$ is a configuration of $G$. We say that $x_{S}(v)$ is the state from $S$ assigned to $v\in V$ by $x_{S}$. We define a task $T=(O,P)$ as a set of states $O$ and a problem specification $P$, where $P$ is a set of configurations over $O$ that we call the correct configurations for $T$. It is possible to have several correct configurations in $P$ over the same graph. The notion of a correct configuration is specific to the task. In this paper, we consider two tasks: leader election (Sections˜3 and 4) and spanning tree construction (Section˜4). We give the definition of correct configurations in the problem specification $P$ and the set of states $O$ for each of these tasks here. We write $x_{O}$ to denote the mapping $x_{O}:V\rightarrow O$.
Leader Election
[…]
Spanning Tree
[…]
  • The set of directed edges $\{\vec{vw}\text{ }|\text{ }v\neq r\}$, where $w=\emph{parent}_{G}(v)$, is a directed spanning tree.
We slightly adapt the definition of local certification of [feuilloley2025certification] to our setting. We assume that the nodes of $G$ do not have identifiers, that is, $G$ is anonymous. Let $L$ be a possibly empty set of input labels. If $L\neq\emptyset$, we say that $G$ is labeled and we call the mapping from $V$ to $L$ the input function. Let $C$ be a non-empty set of states, called the certificate. A certificate assignment on $G$ is a mapping $x_{C}:V\rightarrow C$. For a given certificate assignment $x_{C}$ and a node $v\in V$, the view of $v$ consists of the 1-ball $B_{1}(v)$ and the restriction of the port numbering function $\delta$, the input function (if defined) and the certificate assignment $x_{C}$ to $B_{1}(v)$. An example of the view of a node is given in Figure˜1.
Figure 1: Right: A graph with port labels (red) where each node is assigned a certificate in $\{0,1\}^{*}$ (black). Left: The view of the node marked in gray.
Let $R$ be a set of configurations of the form $(B,x_{C})$ where $B$ is a ball of radius 1 and $x_{C}$ is a certificate assignment defined on the nodes of $B$. We refer to $R$ as a set of local configurations. A local configuration is said to be correct with respect to $(C,R)$ if for every $v\in V$ the pair $(B_{1}(v),x_{C}|_{N[v]})$ belongs to $R$, where $x_{C}|_{N[v]}$ denotes the restriction of $x_{C}$ to $N[v]$. In other words, the local configuration is correct if the labeling in each ball of radius 1 satisfies the local conditions defining a correct configuration $R$. We refer to the pair $(C,R)$ as a certification scheme.
A certification scheme is used to solve a given task $\mathcal{T}=(O,P)$. The key property we prove, is that if a configuration is locally correct with respect to the certification scheme, then it is globally correct (i.e., it is a correct configuration for the task). Formally, let $out:C\rightarrow O$. We say that there exists a certification scheme $(C,R)$ for some task $(O,P)$ in a family of possibly labelled graphs $\mathcal{F}$ if for every $G\in\mathcal{F}$ the following two properties are satisfied [^1] The symbol $\circ$ denotes the composition operator.:
Existence
$\forall G\in\mathcal{F},\text{ }\exists(G,x_{C})\text{ s.t. }\forall v\in V,\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R$
Soundness
$\forall(G,x_{c})\text{ s.t. }G\in\mathcal{F},\text{ if }\forall v\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R\text{ then }(G,\emph{out}\circ x_{C})\in P$
In most of our certification schemes, the certificates correspond to an edge orientation. For these certificates, we denote the configuration as $\vec{G}$ and we adopt the following encoding for the certificates: the certificate assigned to each node $v\in V$ is composed of $\deg(v)$ bits where each bit encodes the direction of the corresponding incident edge. We assume that 0 denotes an incoming edge to $v$ and 1 denotes an outgoing edge from $v$. The bit at position $i\in\{0,\ldots,\deg(v)-1\}$ of the certificate corresponds to the direction of the edge at port $i$ of $v$. A visualization of this for the graph in Figure˜1 is given in Figure˜2.
Figure 2: The $i$-th bit of the certificate (black numbers) of each node $v$ corresponds to the direction of the edge incident to port $i$ (red numbers) of $v$, for $i\in\{0,\ldots,deg(v)-1\}$ where 0 denotes an incoming and 1 an outgoing edge. On the left we show the direction of the edges and on the right we additionally show the bit of the certificate of each node corresponding to each edge.
For each edge $uv$ where $u$ (resp. $v$) is connected to $v$ (resp. $u$) through port $i$ (resp. $j$), we say that the edge is oriented if the $i$-th bit of $u$ differs from the $j$-th bit of $v$.
2.2 Warm Up: A Certificate for Leader Election in Trees
We give a simple certificate that guarantees a unique sink in the class of trees. While not new (e.g., one could easily construct a certificate following the self–stabilizing leader election algorithm of [antonoiu1996trees]), we believe it gives a good intuition behind the certificates and proofs in the following sections. The certificate of each node consists of the orientation of edges incident to the node. A correct configuration is defined by two conditions, T1 and T2. If both T1,T2 hold at every node of a tree $T$, then $T$ has a unique sink which we define to be the leader.
  1. T1
    Each edge is directed.
  2. T2
    For each vertex $v$, $v$ has at most one outgoing edge.
An orientation satisfying $T_{1}$ and $T_{2}$ (i.e., satisfying the Existence property) exists: first direct each edge incident to a leaf toward its unique neighbor, then iteratively orient edges away from leaves in the subgraph of remaining unoriented edges.
Theorem 2.1 (Soundness).
In an unlabeled tree $T$, any configuration $\vec{T}$ satisfying T1 and T2 contains a unique sink.
Proof 2.2.
We prove the theorem by induction on the number of vertices of $T$. The base case of a single vertex is trivial. Suppose $T$ contains at least two vertices and let $v$ be a leaf of $T$. Let $\vec{T}$ be a configuration of $T$ and assume that $\vec{T}$ satisfies T1 and T2. Observe that the configuration $\vec{T}\setminus\{v\}$ of $T\setminus\{v\}$ also satisfies T1 and T2. By the induction hypothesis $\vec{T}\setminus\{v\}$ has a unique sink $t$. We show that $\vec{T}$ also has a unique sink.
Call $u$ the neighbor of $v$ in $T$. Assume $vu$ is oriented $\overrightarrow{uv}$. Then from T2, $u$ has no other outneighbor in $\vec{T}$, so $u$ is the unique sink in $\vec{T}\setminus\{v\}$. Since $u$ has an outgoing edge in $\vec{T}$, it is not a sink, but $v$ is. Thus, $v$ is the unique sink in $\vec{T}$. Assume now $vu$ is oriented $\overrightarrow{vu}$. Then $t$ remains a sink in $\vec{T}$ and $v$ is not a sink, so $t$ is the unique sink. Hence, in all cases $\vec{T}$ has a unique sink, completing the proof.
3 Chordal Graphs
We generalize the certification scheme from Section 2.2 to ensure the validity of a certificate assignment, which guarantees acyclicity and a unique sink in chordal graphs. Each node’s certificate is the orientation of its incident edges. We refer to the nodes of $G$ and the certificate assignment as a configuration $\vec{G}$. A graph is chordal if every induced cycle is of length three. Every chordal graph has a simplicial vertex $v$, that is, a vertex $v$ such that $N(v)$ is a clique. We define a configuration to be correct for leader election, if for every vertex $v$ of a chordal graph, the following local conditions hold:
  1. C1
    Each edge incident to $v$ is directed.
  2. C2
    All outneighbors of $v$ are adjacent, that is, $N^{+}(v)$ induces a complete graph.
  3. C3
    Vertex $v$ does not belong to any directed triangle.
Observe that the verification of C3 is possible, since the view by definition contains the orientation of edges between neighboring nodes.
{observation}
[Existence] A correct configuration with respect to C1–C3 exists and can be constructed as follows: direct all edges incident to a simplicial vertex $u$ toward $N(u)$, then iteratively restrict the graph to the subgraph of remaining unoriented edges and orient edges away from simplicial vertices.
Lemma 3.1.
In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 is an acyclic orientation of $G$.
Proof 3.2.
Suppose the lemma statement does not hold and $\vec{G}$ contains an oriented cycle. Take $\mathsf{C}=(u_{0},u_{1},\ldots,u_{k-1})$ to be the smallest such cycle. From C3, $k>3$. Then without loss of generality, $u_{1}$ is adjacent to at least three edges: $\overrightarrow{u_{0}u_{1}}$, $\overrightarrow{u_{1}u_{2}}$ and some directed edge $u_{1}u_{i}$ for $3\leq i<k$. Depending on the orientation of $u_{1}u_{i}$, $G$ contains either the oriented cycle $\mathsf{C}^{\prime}=(u_{1},u_{i},u_{i+1},\ldots,u_{k-1},u_{0})$ or the oriented cycle $\mathsf{C}^{\prime\prime}=(u_{1},u_{2},\ldots,u_{i},u_{1})$. In either case, a smaller cycle than $\mathsf{C}$ exists in $G$, contradiction.
Lemma 3.3 (Soundness).
In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 contains a unique sink.
Proof 3.4.
We prove the lemma by induction on the number of vertices of $G$. The base case holds for $G$ having one vertex. For $G$ with at least two vertices, let $u$ be a simplicial vertex, that is, $N(u)$ induces a complete subgraph. Consider a configuration $\vec{G}$ of $G$ satisfying C1–C3. Removing $u$ from $\vec{G}$ gives a configuration of $G\setminus\{u\}$ that also satisfies C1–C3, so by the induction hypothesis $\vec{G}\setminus\{u\}$ has a unique sink $t$.
We claim that if $u$ has an outneighbor $v$ in $\vec{G}$, then $t$ is also the unique sink in $\vec{G}$. Due to $\vec{G}\setminus\{u\}$ having a unique sink, no vertex other than $t$ can be a sink in $\vec{G}$. If $t$ is not a sink in $\vec{G}$, then $\overrightarrow{tu}\in\vec{G}$, as all remaining edges incident to $t$ are incoming due to $t$ being a sink in $\vec{G}\setminus\{u\}$. Since $v,t\in N(u)$ and $u$ is a simplicial vertex, $v\in N^{-}(t)$. Thus $\overrightarrow{tu},\overrightarrow{uv},\overrightarrow{vt}$ induce a directed triangle, contradicting C3.
Suppose now that $u$ has no outneighbor, (i.e., $u$ is a sink of $\vec{G}$). For any neighbor $v$ of $u$, any outneighbor $w$ of $v$ in $\vec{G}\setminus\{u\}$ must be a neighbor of $u$ in $\vec{G}$ from C2. Consequently, $N^{+}(v)\subseteq N[u]$. We claim that there exists a vertex $t^{\prime}\in N(u)$ that has no outneighbor in $N(u)$. Suppose otherwise. Then, every $v\in N(u)$ has an outneighbor in $N(u)$ and $\vec{G}\setminus\{u\}$ contains a cycle, which is a contradiction by Lemma 3.1. Since $N(u)$ is a complete graph, this $t^{\prime}$ is unique. Furthermore, as we have already shown that $t^{\prime}$ has no outneighbor outside of $N[u]$, $t^{\prime}$ is the unique sink of $\vec{G}\setminus\{u\}$. Since $^{\prime}\in N^{-}(u), tt$ is not a sink in $\vec{G}$. Therefore, $u$ is the unique sink of $\vec{G}$.
Theorem 3.5 directly follows from Section˜3 and Lemmas˜3.1 and 3.3.
Theorem 3.5.
In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 is an acyclic orientation of $G$ containing a unique sink.
4 Dismantlable graphs
This section is divided into two parts: a certification scheme ensuring a unique sink in $K_{4}$-free dismantlable graphs and a certification scheme for a rooted spanning tree, given a unique root, in dismantlable graphs. Let $G$ be a graph and $u,v$ be two vertices in $V(G)$. We say that $v$ is dominated by $u$ in $G$ if $N[v]\subseteq N[u]$. A graph is dismantlable if there exists an ordering $v_{1},\ldots,v_{n}$ of $V(G)$ such that in $G[\{v_{1},\ldots,v_{i}\}]$, $v_{i}$ is dominated by some $v_{j}$ with $j<i$. So for each $i>1$, there exists $j<i$ such that $N[v_{i}]\cap\{v_{1},\ldots,v_{i}\}\subseteq N[v_{j}]$. Dismantlable graphs are exactly the class of cop-win graphs in the classic Cops and Robbers game [aigner1984cops, nowakowski1983vertex, quilliot1978jeux] and include well-structured families such as Bridged graphs and Helly graphs, which are known for their convexity and metric properties.
4.1 Leader election in $K_{4}$-free dismantlable graphs
In this section we only consider $K_{4}$-free dismantlable graphs. As in Section 3, the certificate at each node consists of an orientation on the edges incident to the node and we write $\vec{G}$ to denote a configuration. We define a locally correct configuration to be one that satisfies the following rules for every node $v$ of $G$. We show that if all rules are satisfied, then $G$ contains a unique sink that we define to be the leader.
  1. D1
    Each edge incident to $v$ is directed.
  2. D2
    Either $v$ has no outgoing edges, or there exists an outneighbor $v^{+}$ of $v$ adjacent to all other outneighbors of $v$, i.e, $N^{+}(v)\subseteq N[v^{+}]$.
  3. D3
    Vertex $v$ does not belong to any directed triangle.
{observation}
[Existence] Any dismantlable graph admits an acyclic orientation satisfying D1–D3. {claimproof} Let $v_{1},v_{2}\ldots,v_{n}$ be the dismantling order of the nodes of some $K_{4}$-free dismantlable graph $G$. Orient each edge between two nodes $v_{i}$ and $v_{j}$ as $\overrightarrow{v_{i}v_{j}}$ if and only if $i>j$ in the dismantling order. This orientation satisfies D1 and D3. For D2, consider any node $v$. If $N^{+}(v)=\emptyset$, D2 is immediately satisfied for $v$. So suppose $N^{+}(v)\neq\emptyset$. By dismantlability, every node $v$ is dominated by some later node $u$ in the dismantling order. Then $\overrightarrow{vu}$ is an edge of the orientation and for every other outneighbor $w$ of $v$, $w\in N(u)$. Hence, $N^{+}(v)\subseteq N[u]$, satisfying D2.
However, notice that in the case of $K_{4}$-free dismantlable graphs, one cannot hope to have a property similar to Lemma˜3.1. Indeed, as in [chalopin2024stabilising], one can construct configurations satisfying D1-D3 that contain directed cycles. An example is shown in Figure˜3.
Figure 3: A $K_{4}$-free dismantlable graph and a certificate assignment that satisfies D1–D3 but is not acyclic. The square node dominates all cyclic nodes. The edges forming the cycle are marked in orange.
Theorem 4.1 (Soundness).
Let $G$ be a $K_{4}$-free dismantlable graph. If a configuration $\vec{G}$ satisfies D1–D3, then $\vec{G}$ contains a unique sink.
Proof 4.2.
Consider a graph $G$ and a configuration $\vec{G}$ of $G$ satisfying conditions D1–D3. We prove the theorem by induction on the number of vertices in $G$. The theorem trivially holds if $G$ has one vertex. Suppose $G$ has at least two vertices. Let $v$ be a vertex dominated by some vertex $u$ in $G$, and let $G^{\prime}=G\setminus\{v\}$ and $\vec{G}^{\prime}=\vec{G}\setminus\{v\}$. Since D1 and D3 hold in $\vec{G}$, they also hold in $\vec{G}^{\prime}$. To apply the induction hypothesis to $\vec{G}^{\prime}$, it remains to show that D2 holds in $\vec{G}^{\prime}$. For any vertex $w$ such that $v\notin N^{+}(w)$ in $\vec{G}$, condition D2 remains satisfied for $w$ in $\vec{G}^{\prime}$. More precisely, if D2 is not satisfied for $w$ in $\vec{G}^{\prime}$, then $v\in N^{+}(w)$ and $v$ is the only possible choice for $w^{+}$ in $\vec{G}$. We first establish a useful lemma.
Lemma 4.3.
For any vertex $w\in N^{-}(v)\cap N^{-}(u)$, we can assume that $w^{+}\neq v$.
Proof 4.4.
If $w^{+}=v$, i.e., $N^{+}(w)\subseteq N[v]$, then $N^{+}(w)\subseteq N[v]\subseteq N[u]$ and we can thus choose $w^{+}=u$.
In the remainder of the proof we show that D2 is satisfied in $\vec{G}^{\prime}$ and that $\vec{G}$ contains a unique sink. We distinguish two cases, depending on whether $uv$ is directed as $\overrightarrow{uv}$ or $\overrightarrow{vu}$.
Case 1.
The edge $uv$ is $\overrightarrow{vu}$.
{claimproof}
Consider a vertex $w\neq v$ such that $w\in N(v)$. Observe that if $w\in N^{+}(v)$ and D2 is satisfied in $\vec{G}$ for $w$ then D2 cannot be broken for $w$ in $\vec{G}^{\prime}$. So suppose that $w\in N^{-}(v)$. Since $u\in N^{+}(v)$, by D3, $w\in N^{-}(u)$. Consequently, by Lemma˜4.3, we can assume that $w^{+}\neq v$ and D2 is thus satisfied for $w$ in $\vec{G}^{\prime}$.
Since D1–D3 are satisfied in $\vec{G}^{\prime}$, $\vec{G}^{\prime}$ contains a unique sink $t$. Observe that if $t\in N^{-}(v)$ in $\vec{G}$, $\overrightarrow{tv},\overrightarrow{vu},\overrightarrow{ut}$ induce a directed cycle in $\vec{G}$, a contradiction. Consequently, $t$ is also a sink in $\vec{G}$. Since all other vertices of $\vec{G}^{\prime}$ have outgoing edges in $\vec{G}^{\prime}$ (and thus in $\vec{G}$), and since $v$ is not a sink in $\vec{G}$ (as $v\in N^{-}(u)$), $t$ is the unique sink of $\vec{G}$. This ends the proof of ˜1.
Case 2.
The edge $uv$ is $\overrightarrow{uv}$.
{claimproof}
We further distinguish two cases, depending on whether we can assume that $u^{+}=v$.
Case 2.1.
$N^{+}(u)\subseteq N[v]$, that is, we can assume $u^{+}=v$.
{claimproof}
Since $v$ is dominated by $u$, for any $w\in N(v)$ we also have $w\in N(u)$. If additionally $w\in N^{-}(u)$, then by D3 it follows that $w\in N^{-}(v)$. Consequently, for any $w\in N(v)$, it must be that $w\in N^{+}(u)\cap N^{+}(v)$, or $w\in N^{-}(u)\cap N^{-}(v)$, or $w\in N^{+}(u)\cap N^{-}(v)$.
We distinguish two cases, depending on whether $N^{+}(u)\cap N^{-}(v)$ is empty.
Case 2.1.1.
$N^{+}(u)\cap N^{-}(v)=\emptyset$
{claimproof}
We first show that D2 is satisfied in $\vec{G}^{\prime}$. Consider a vertex $w\neq v$ such that $w\in N^{-}(v)$. If $w\neq u$, since $u$ dominates $v$, $u$ must be a neighbor of $w$. Given that $N^{+}(u)\cap N^{-}(v)=\emptyset$, it follows that $w\in N^{-}(u)$. By Lemma˜4.3, we can assume $w^{+}\neq v$, thus D2 is satisfied for $w$ in $\vec{G}^{\prime}$.
We now show that D2 holds for $u$. Since $N^{+}(u)\subseteq N[v]$ and $N^{+}(u)\cap N^{-}(v)=\emptyset$, we have $N^{+}(u)\subseteq N^{+}(v)\cup\{v\}$. Given that $N^{+}(v)\subseteq N(v)\subseteq N[u]$ and $v\in N^{+}(u)$, by D3, we obtain $N^{+}(v)\subseteq N^{+}(u)$. Therefore, $N^{+}(u)=N^{+}(v)\cup\{v\}$. Hence, either $N^{+}(v)=\emptyset$ and $N^{+}(u)=\{v\}$, or there exists $v^{+}\in N^{+}(v)$ such that $N^{+}(v)\subseteq N[v^{+}]$, implying $N^{+}(u)=N^{+}(v)\cup\{v\}\subseteq N[v^{+}]$. In either case, D2 is satisfied for $u$ in $\vec{G}^{\prime}$.
By induction hypothesis $\vec{G}^{\prime}$ contains a unique sink $t$. Notice that as $N^{+}(u)=N^{+}[v]$, $v$ is a sink of $\vec{G}$ if and only if $u$ is a sink of $\vec{G}^{\prime}$. Therefore, if $t\neq u$ in $\vec{G}^{\prime}$, $v$ is not a sink in $\vec{G}$ and $t$ is the unique sink in $\vec{G}$. If $t=u$, then $t$ is not a sink in $\vec{G}$ but $v$ is, consequently $v$ is the unique sink of $\vec{G}^{\prime}$. In either case, $\vec{G}$ contains a unique sink concluding the proof of ˜2.1.1.
Case 2.1.2.
$N^{+}(u)\cap N^{-}(v)\neq\emptyset$
{claimproof}
In this case, we consider the configuration $\vec{G}^{*}$ obtained from $\vec{G}$ where for each $w\in N^{+}(u)\cap N^{-}(v)$ in $\vec{G}$, we have replaced the arc $\overrightarrow{uw}$ by the arc $\overrightarrow{wu}$. We show that D1–D3 are satisfied in $\vec{G}^{*}$. Observe that since all edges are directed in $\vec{G}$, they are all directed in $\vec{G}^{*}$ and D1 holds.
Suppose that there exists a directed triangle in $\vec{G}^{*}$. Then necessarily, this triangle contains an edge $\overrightarrow{wu}$ with $w\in N^{+}(u)\cap N^{-}(v)$ in $\vec{G}$. Let $x$ be the vertex such that $\overrightarrow{wu}$, $\overrightarrow{ux}$, $\overrightarrow{xw}$ form a triangle in $\vec{G}^{*}$. Note that $x\neq v$ as $w\in N^{-}(v)\cap N^{+}(x)$ in $\vec{G}$ and $\vec{G}^{*}$. Since $\overrightarrow{ux}$ is in $\vec{G}^{*}$, $x\in N^{+}(u)$ in $\vec{G}^{*}$ and in $\vec{G}$ (the set of outneighbors of $u$ in $\vec{G}^{*}$ is a subset of its set of outneighbors in $\vec{G}$). Since in $\vec{G}$, $N^{+}(u)\subseteq N[v]$, $x$ is a neighbor of $v$. Since $w\in N^{+}(x)$ and $v\in N^{+}(w)$, by D3, we have $x\in N^{-}(v)$ in $\vec{G}$. Consequently, in $\vec{G}$, $x\in N^{+}(u)\cap N^{-}(v)$, and thus in $\vec{G}^{*}$, $x\in N^{-}(u)$, a contradiction. This shows that D3 holds in $\vec{G}^{*}$.
We now show that D2 holds in $\vec{G}^{*}$. For any $\notin(N^{+}(u)\cap N^{-}(v))\cup\{u\}, xx$ has the same set of outneighbors in $\vec{G}$ and in $\vec{G}^{*}$, and thus D2 holds for $x$ in $\vec{G}^{*}$ as it holds in $\vec{G}$. Since $N^{+}(u)\subseteq N(v)$ in $\vec{G}$, since $N^{+}(u)$ in $\vec{G}^{*}$ is a subset of $N^{+}(u)$ in $\vec{G}$, and since $v\in N^{+}(u)$ in $\vec{G}^{*}$, D2 also holds for $u$ in $\vec{G}^{*}$ (with $u^{+}=v$). Consider now $w\in(N^{+}(u)\cap N^{-}(v))$ and let $w^{+}\in N^{+}(w)$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $\vec{G}^{\prime}$. As $u\notin N^{+}(w)$ in $\vec{G}$, $w^{+}\neq u$. Since $v\in N^{+}(w)$, $w^{+}\in N[v]\setminus\{u\}\subseteq N(u)$. Hence, $w^{+}\in N(u)$, implying $u\in N[w^{+}]$ and thus in $\vec{G}^{*}$ we also have $N^{+}(w)\subseteq N[w^{+}]$. Therefore, D2 is satisfied at $w$ in $\vec{G}^{*}$. Hence D2 holds in $\vec{G}^{*}$.
Consequently, by ˜2.1.1 applied to $\vec{G}^{*}$, we deduce that there exists a unique sink $t$ in $\vec{G}^{*}$. Note that $t\notin(N^{+}(u)\cap N^{-}(v))\cup\{u\}$ as every vertex $w\in(N^{+}(u)\cap N^{-}(v))\cup\{u\}$ has $v$ as an outneighbor in $\vec{G}$ and in $\vec{G}^{*}$. So, $t$ is a sink in $\vec{G}$ and since no node of $(N^{+}(u)\cap N^{-}(v))\cup\{u\}$ is a sink in $\vec{G}$, $t$ is the unique sink of $\vec{G}$. This ends the proof of ˜2.1.2. This ends the proof of ˜2.1.
Case 2.2.
$N^{+}(u)\not\subset N[v]$, implying that $u^{+}\neq v$.
{claimproof}
We first show that D2 is satisfied in $\vec{G}^{\prime}$. For D2 to be violated, $v$ should be the only choice for $w^{+}$, for some $w\neq u$. If $N^{+}(w)=\{v\}$, D2 trivially holds for $w$ in $\vec{G}^{\prime}$. If $u\in N^{+}(w)$, D2 holds for $w$ by Lemma˜4.3. Suppose now there exists $z\in N^{+}(w)\setminus\{v\}$. Then $z\in N(v)$ and from $u$ dominating $v$, $\{w,z\}\in N(u)$. Hence, $v,u,w,z$ induce a $K_{4}$, contradiction. So we can assume that $w^{+}\neq v$ for all $w\in V$, and D2 holds for all nodes in $\vec{G}^{\prime}$.
By induction hypothesis, $\vec{G}^{\prime}$ contains a unique sink $t$. We consider two cases, based on the orientation of $u^{+}v$.
Case 2.2.1.
The edge $u^{+}v$ is $\overrightarrow{vu^{+}}$, that is, $u^{+}\in N^{+}(v)$.
{claimproof}
If $t$ is not a sink in $\vec{G}$, $t\in N^{-}(v)$ and thus, $t\in N(u)$. Observe that it cannot be that $t=u^{+}$ since we have assumed $\overrightarrow{vu^{+}}$, which means that $vt$ is oriented as $\overrightarrow{vt}$ and $t$ would still be a sink in $\vec{G}$ contradicting our assumption. So suppose $t\neq u^{+}$. Since $t$ is a sink in $\vec{G}^{\prime}$, $t\in N^{+}(u)$ and thus $t\in N(u^{+})$, forming a $K_{4}$ among $u,v,t,u^{+}$. Therefore, $t\notin N^{-}(v)$ and $t$ is also a sink in $\vec{G}$. Since $v$ is not a sink in $\vec{G}$ due to $\overrightarrow{vu^{+}}$, $t$ is the unique sink in $\vec{G}$. This ends the proof of ˜2.2.1.
Case 2.2.2.
The edge $u^{+}v$ is $\overrightarrow{u^{+}v}$, that is, $u^{+}\in N^{-}(v)$.
{claimproof}
We claim that $u^{+}=t$. Suppose for contradiction that $u^{+}\neq t$. Then, since D2 is satisfied in $\vec{G}$, there exists $u^{++}\in N^{+}(u^{+})$ such that $N^{+}(u^{+})\subseteq N[u^{++}]$. Since we have assumed that $v\in N^{+}(u^{+})$, it must be that $v\in N[u^{++}]$. Suppose $v\neq u^{++}$, that is, $v\in N(u^{++})$. Then since $v$ is dominated by $u$, the node $u^{++}\in N(u)$. However, this means that the vertices $v,u,u^{+}$ and $u^{++}$ induce a $K_{4}$, hence this case is not possible. So suppose $v=u^{++}$. Suppose that $u^{+}$ has some outneighbor $w$ distinct from $v$. Then, since $v=u^{++}$, $w$ must be in $N(v)$ and $u$ is connected to $w$ as $u$ dominates $v$. In this case, the vertices $v,u,u^{+}$ and $w$ induce a $K_{4}$ so this setting is not possible and it must be that $N^{+}(u^{+})=\{u^{++}\}=\{v\}$. Therefore, it must be that $u^{+}=t$ in $\vec{G}^{\prime}$, contradicting the initial assumption.
Consider now $u^{+}=t$ in $\vec{G}^{\prime}$, implying that $v$ is the only outneighbor of $u^{+}$ in $\vec{G}$. Let $z\in N(v)$. Then, $z$ must also be neighboring to $u$ as $u$ dominates $v$. We claim that $v$ is the unique sink in $\vec{G}$. The orientation $\overrightarrow{uz}$ means that the edge $u^{+}z$ must also exist implying that $u,v,u^{+}$ and $z$ induce a $K_{4}$ which is not possible. So $uz$ must be oriented as $\overrightarrow{zu}$. From D3, $zv$ should be oriented as $\overrightarrow{zv}$. Hence every neighbor of $v$ in $\vec{G}$ must be incoming to $v$. Therefore, $v$ is a sink in $\vec{G}$ and since the unique sink $u^{+}$ of $\vec{G}^{\prime}$ is not a sink in $\vec{G}$ due to $\overrightarrow{u^{+}v}$, $v$ is the unique sink in $\vec{G}$. This ends the proof of ˜2.2.2. This ends the proof of ˜2.2. This ends the proof of ˜2. Theorem 4.1 follows from the analysis in Cases 1 and 2.
4.2 Spanning Tree
We present a certificate verifying the existence of a spanning tree in any dismantlable labeled graph $G=(V,E)$. The labeling function assigns each $v\in V$ a label in $L=\{0,1\}$, with exactly one node labeled 1 (the root $r$) and all others labeled 0. We define $C=\{0,1,2\}$ to be the certificate and each node is assigned a value from $C$ by $x_{C}:V\rightarrow C$. The following must hold for every node $v$ of $G$ with value $i$.
  1. N1
    Node $v\neq r$ has at least one neighbor with value $(i-1)\mod 3$.
  2. N2
    Node $v$ is not in any three-node triangle labeled $\{0,1,2\}$.
We show that if N1–N2 are satisfied for every node in $G$, adding an edge from any node $v$ distinct from the root with value $i$ to any neighbor of $v$ with value $i-1\mod 3$, induces a spanning tree.
Lemma 4.5 (Existence).
In any graph $G$ there exists a certificate assignment that satisfies N1–N2 for every node in $G$.
Proof 4.6.
We write $d(v,r)$ to denote the distance between a node $v$ and the root $r$. We construct the assignment as follows. For each node $v$ at distance $k>1$ from the root in $G$, assign $d(v,r)\mod 3$ to $v$. This satisfies N1. Let $u,w$ be two neighbors of $v$ that are also neighboring to each other. Without loss of generality assume that $v$ is the closest vertex to $r$ among the triangle $u,v,w$ and let $k=d(v,r)$. Then $k\leq d(u,r)\leq k+1$ and $k\leq d(w,r)\leq k+1$, thus N2 is satisfied and an assignment that satisfies N1–N2 exists.
[…]
Lemma 4.7.
If N1–N2 are satisfied for every node in a dismantlable graph $G$, $\vec{G}$ does not contain oriented cycles.
Proof 4.8.
Suppose the lemma does not hold. We define the order of a vertex $v$ as its position in the dismantling order. Let the sum of orders of a cycle be the sum of the orders of its vertices. Among all directed cycles in $\vec{G}$, let $\mathsf{C}=(v_{1},v_{2},\ldots,v_{k})$ be the cycle that minimizes this sum. Notice that the values of all nodes in $\mathsf{C}$ must follow the sequence $\ldots,0,1,2,0,1,2,\ldots$ so that the orientation of the cycle remains consistent. Call $v$ the node with the largest order in $\mathsf{C}$, $u$ the node dominating $v$ and let $w,z$ be the two neighbors of $v$ in $\mathsf{C}$. Observe that $u$ is also adjacent to $w,z$ and that by definition its order is smaller than the one of $v$. Without loss of generality let the value of $v$ be 1 and let the values of $w,z$ be 0 and 2 respectively. Then, the value of $u$ must be 1, otherwise one of $(u,v,w)$ and $(u,v,z)$ breaks N2. We show that there exists a cycle $\mathsf{C}^{\prime}$ with a smaller sum of orders than $\mathsf{C}$ which includes $u$ instead of $v$. First suppose that $u\in\mathsf{C}$ and let $u_{-}$ be the predecessor of $u$ in $\mathsf{C}$. Then $\mathsf{C}^{\prime}$ is induced by the nodes in $\mathsf{C}\setminus\{v,z,\ldots,u_{-}\}$. Suppose now that $u\notin\mathsf{C}$ and let $\mathsf{C}^{\prime}$ be the cycle induced by the nodes in $\{u\}\cup\mathsf{C}\setminus\{v\}$. In either case, the existence of $\mathsf{C}^{\prime}$ contradicts $\mathsf{C}$ being a cycle with minimum sum of orders, proving the lemma.
Let $T$ be a spanning subgraph of $\vec{G}$ in which each node $v$, except the root, arbitrarily selects a parent among its neighbors with value equal to $(\textit{value of }v-1)\mod 3$. We show that $T$ is a tree. Since $\vec{G}$ is acyclic by Lemma 4.7, it follows that $T$ is also acyclic. We prove that $T$ is connected.
Lemma 4.9.
If N1–N2 are satisfied for every node in a dismantlable graph, $T$ is connected.
Proof 4.10.
Let $u$ be a node in $T$. Since from Lemma 4.7, $\vec{G}$ and as a result $T$ does not contain cycles, the directed path starting at $u$ must end at a node without a successor in $T$. From N1, the root is the only node without a smaller neighbor. Hence, any node $u$ must be in a path that reaches the root, and as a result $T$ is connected.
Finally, since every node in $G$ has a label and must be in $T$, from Lemmas 4.7 and 4.9 we obtain the following:
Theorem 4.11 (Soundness).
In a dismantlable graph, if N1–N2 are satisfied for every node, $T$ is a spanning tree.
5 From Certificate to Self–Stabilization
An algorithm is called self-stabilizing when it can autonomously recover from transient faults. In particular, a distributed self-stabilizing algorithm is usually defined as a set of rules evaluated at every node. Each rule is of the form:
$\texttt{name-of-rule}:\langle guard\rangle\longrightarrow command$
where guard is a boolean predicate over the variables of the closed neighborhood of the node and the command specifies how the node updates its variables. Here, we assume that each node only updates its state variable. A self-stabilizing algorithm is said to be silent if after some finite time no node updates its variables anymore. When activated, each node can perform computations based on its state and the states of its neighbors and update its own state. The order of node activations is determined by a scheduler. A scheduler is characterized both by synchronicity and fairness. In particular, here we consider an asynchronous scheduler meaning that at each step, an arbitrary (non-empty) subset of nodes may be activated. In terms of fairness, we assume a Gouda fair scheduler [Gouda2001theory]. According to [dubois2011taxonomy], Gouda fairness is the strongest fairness assumption. It requires that for any configuration that appears infinitely often during the execution of an algorithm, every configuration that is reachable by executing any non-empty subset of enabled commands (i.e., commands for which the guard is true) also appears infinitely often. We assume that the graphs we consider are finite. In this setting, Gouda fairness ensures that any configuration that is infinitely often reachable is eventually reached. In this section we show how to obtain a silent self–stabilizing algorithm that works under a Gouda fair scheduler, given a certification scheme.
We say that under a given algorithm $A$, a certificate assignment is final if no node satisfies any of the guards in $A$. We define a correct final certificate assignment to be one that is both final with respect to an algorithm $A$ and correct with respect to a certification scheme. Call the set of states in a given certificate $C$. Informally, every node that detects that a condition of the certification scheme is violated sets its state to a reset state that we denote $\bot$. Additionally, every node that has at least one neighbor that has detected an error and reset its state, also changes its state to the reset state. Hence, an error detection is propagated throughout the graph from node to neighboring node. When a node that is at the reset state is activated, it non-deterministically switches to a state from $C$. Gouda fairness ensures that the graph eventually reaches a correct global configuration.
A node executing Algorithm 1 evaluates the guards in order. If the first guard is enabled, the node performs the corresponding command and deactivates without evaluating the second guard.
Input: The certification scheme $(C,R)$ and a certificate assignment $x_{C}:V\rightarrow C$
1exEvery time node $v$ is activated:
/* states from $C$ are always chosen non-deterministically */
S1 : v.state $\neq\bot$ $\land$ $((B_{1}(v),x_{C}|_{N[v]})\notin R$ $\lor$ $\exists u\in N(v):$ u.state = $\bot)$ $\rightarrow$ v.state $\leftarrow\bot$
S2 : v.state = $\bot$ $\rightarrow$ v.state $\leftarrow$ state from $C$
Algorithm 1 General Certificate to Self-Stabilizing Algorithm
Theorem 5.1.
Let $(C,R)$ be a certification scheme satisfying the Existence and Soundness conditins for some task $\mathcal{T}$ in a connected graph $G$. Starting from any arbitrary configuration, any Gouda fair execution of Algorithm 1 on $G$ stabilizes to a correct final configuration.
Proof 5.2.
Let $G^{*}$ be a configuration where no node satisfies the guard of S1 or of S2, that is, $G^{*}$ is a final correct configuration. By definition, $\mathcal{T}$ is solved in $G^{*}$, since $(C,R)$ is a certification scheme for $\mathcal{T}$. Suppose for contradiction that there exists a configuration $G^{\prime}\neq G^{*}$ that is infinitely often reachable during the execution of Algorithm 1. Since $G^{*}$ is final, it cannot be reachable from $G^{\prime}$. As $G^{\prime}$ is not final, either $G^{\prime}=G_{\bot}$ (i.e., all nodes are in state $\bot$) or the guard of S1 is enabled for some node. In the former case, $G_{\bot}$ trivially appears infinitely often. In the latter, by S1, there exists an activation sequence that leads to $G_{\bot}$. By Gouda fairness, for every a configuration that appears infinitely often, every reachable configuration also appears infinitely often. Hence, in either case $G_{\bot}$ appears infinitely often. Due to S2, once all nodes in $G_{\bot}$ are activated, the configuration becomes $G^{*}$, contradiction. Thus, no configuration $G^{\prime}\neq G^{*}$ can appear infinitely often.
A self–stabilizing algorithm can be obtained by combining the respective certification scheme of each of the previous sections with Algorithm 1.
6 Discussion and Open Problems
In Sections˜3 and 4 we gave certification schemes for leader election, with constant size per edge. A natural direction for further research is studying whether our certification scheme from Section˜4 also works for dismantlable graphs. In particular, only the proof of ˜2.2 depends on the graph being $K_{4}$-free. However, removing this restriction has been very challenging. If this is not possible, another direction is considering bridged or Helly graphs, which are dismantlable but enjoy interesting properties that could be useful for addressing our problem. Alternatively, if our certification scheme can be proven to also work on dismantlable graphs, it is worth exploring other relevant classes. For example, [fieux2020hierarchy] gives a hierarchy on classes (based also on the work of [mazurkiewicz2007locally]) that have dismantling orderings, called s-dismantlable, where 1-dismantlable graphs correspond to the dismantlable graphs defined in Section 4. Informally, for $s>1$, [fieux2020hierarchy] defines a node to be $s+1$ dismantlable if the graph induced by its open neighborhood is $s$-dismantlable and a graph to be $s$-dismantlable if it is reducible to a single vertex by successive deletions of $s$-dismantlable vertices. These directions for future research are summarized in Figure˜4.
Figure 4: A graph class hierarchy, where each arrow represents an inclusion relation. The classes highlighted in green are an easy extension of [chalopin2024stabilising] and those in yellow are considered in this paper for leader election certification with constant memory per edge.
Another potential direction for future research is exploring whether it is possible to automatically transform a certificate into a self-stabilizing algorithm with a weaker scheduler, as defined in the taxonomy of [dubois2011taxonomy], even using non-constant memory. For example, the transformation from [BlinFP14], which uses logarithmic memory but works even with an unfair scheduler, contrasts with our algorithm, which uses constant memory but requires a Gouda fair scheduler. These two results suggest that there may be a trade-off between the strength of the scheduler and the memory requirements that has not yet been fully explored.

### v2

@@
-Constant-Size Certificates for Leader Election in Chordal Graphs and Related Classes
+Leader Election via Unique Sink Orientation
 Abstract
-In distributed computing a certification scheme consists of a set of states and conditions over those states that enable each node of a graph to efficiently verify the correctness of a solution to a given problem.
-This work focuses on two fundamental problems: leader election and spanning tree construction.
-For each problem, we present a constant-size (per edge), local certification scheme, where the conditions available to each node can only refer to the graph induced by its one-hop neighborhood.
-In particular, we provide certification schemes for leader election in chordal and $K_{4}$-free dismantlable graphs and for spanning tree construction in dismantlable graphs, assuming a root is given.
-For chordal graphs, our leader election certification scheme additionally ensures an acyclic orientation, a property that is not generally verifiable using constant-size certificates in arbitrary graphs.
-To the best of our knowledge, these are the first local certification results tailored to these graph classes, potentially highlighting structural properties useful for verifying additional problems.
-Finally, we propose an algorithm that automatically transforms any certification scheme into a silent self-stabilizing algorithm (i.e., an algorithm that automatically recovers from faults) by adding only one extra state to the set of states of the certification scheme, assuming a Gouda fair scheduler.
+A Locally Checkable Labeling (LCL) is a specification describing a set of labels that are valid with respect to a set of conditions that characterize a local part of a solution to a global problem.
+Conditions can only refer to nodes and labels within a constant radius neighborhood of each node.
+This work studies local labeling schemes whose global consistency implies solutions to two classical problems: leader election and spanning tree construction.
+For each problem, we present a local labeling scheme using one bit per edge or equivalently $O(\Delta)$ bits per node (where $\Delta$ is the maximum degree in the graph), with conditions checkable within the graph induced by the one neighborhood of each node.
+For leader election, we show that global satisfaction of the conditions implies the existence of a unique sink in the graph, which we define to be a leader, while in the spanning tree setting it implies that a specific subset of edges induces a spanning tree rooted at a given node.
+We show those implications for $K_{4}$-free dismantlable and chordal graphs in the former case and for dismantlable graphs in the latter, assuming a root is given.
+For chordal graphs, the labeling implying a unique sink additionally induces an acyclic orientation.
+This property is not generally locally verifiable with constant-size labels in arbitrary graphs.
+To the best of our knowledge, these are the first local labeling results tailored to ($K_{4}$-free) dismantlable graphs, potentially highlighting structural properties useful for designing LCLs for additional problems.
+Finally, we present a generic transformation that converts any local labeling scheme into a silent self-stabilizing algorithm by adding only one extra state, assuming a Gouda fair scheduler.
 This transformation may be of independent interest.
 keywords:
-Local Certification, Distributed Verification, Proof Labelling Schemes, Chordal Graphs, Dismantlable Graphs, Leader Election, Distributed Computing
+Locally Checkable Labeling, LCL, Leader Election, Dismantlable Graphs, Chordal Graphs, Self-Stabilization, Distributed Computing
 1 Introduction
-Local certification [feuilloley2019introduction], also known as proof labelling schemes [korman2005proof], locally checkable proofs [goos2016locally] or distributed verification, verifies global properties through local conditions.
-Each node individually assesses these conditions and either accepts (i.e., all conditions are satisfied) or rejects (i.e., at least one condition is violated).
-An error is detected if any node rejects.
-For example, to verify that a graph is a cycle and not a path, each node accepts if its degree is two and rejects if its degree is one.
-However, the inverse problem of accepting a path and rejecting a cycle is not possible with certificates of constant size, even if the nodes are not identical [feuilloley2019introduction].
+A fundamental notion in distributed computing is locality, which studies the extent to which global problems admit solutions based solely on local information.
+In a local algorithm, each node in a network must determine its part of a global solution based solely on the topology and states of its radius $r$ neighborhood, for some $r$ that is smaller than the size of the network.
+The challenge of such local algorithms is ensuring that these independently computed local solutions remain globally consistent and, as a result, solve a global task without the benefit of global visibility.
+Locality is concerned with two notions:
+  1. 1.
+    Solving a global problem: If a set of local conditions are satisfied then a global property holds.
+  2. 2.
+    Verifying the solution to a global problem: If a global property holds then a set of local conditions are satisfied.
 The ability to continue functioning and recovering from faults is essential for distributed systems.
-Dijkstra [dijkstra1974stabilising] introduced the concept of self-stabilization as a generic way to model various faults and guarantee the correct behavior of an algorithm regardless of transient faults.
+Dijkstra introduced the concept of self-stabilization [dijkstra1974stabilising] as a generic way to model various faults and guarantee the correct behavior of an algorithm regardless of transient faults.
 An algorithm is self-stabilizing if it eventually converges to a correct output from any arbitrary initial configuration (see also [dolev2000book]).
-A self-stabilizing algorithm is called silent if when it reaches a correct configuration, it remains in that configuration unless a fault occurs.
+A self-stabilizing algorithm is called silent [dolev1999silent] if when it reaches a correct configuration, it remains in that configuration unless a fault occurs.
 On the other hand, a self-stabilizing algorithm that is not silent can transition between correct configurations as the algorithm perpetually continues computing.
 The design of a self-stabilizing algorithm typically consists of three components: (a) Detect when the global configuration is faulty (b) Reach a non-faulty configuration (c) Solve the original task.
-The first aspect naturally connects self-stabilization with local certification, where nodes maintain local certificates that allow verification of the global configuration via local checks.
-For instance, in the self-stabilizing spanning tree algorithm of [chen1991spanning], each node stores a pointer to its parent (as a way to solve the task) and its distance from the root (as a way to verify the solution).
+Thus self-stabilization is naturally connected to the concept of locality as observed for the first time in [afek1997local].
+In a local algorithm each node maintains: (i) a label which is the local part of a global solution and (ii) a certificate which is additional information that allows the verification of the global configuration via local checks.
+For instance, in the self-stabilizing spanning tree algorithm of [chen1991spanning], each node stores a label which is a pointer to its parent (as a way to solve the task) and a certificate which is its distance from the root (as a way to verify the solution).
 Without the latter, cycles cannot be locally detected.
-As a form of fault-tolerance, verification is not the main task to be solved within a system and as a result it should not consume too many resources.
-Hence, in this work, we only consider certificates of constant size per edge.
-A significant challenge in this setting is that in general problems such as the detection of cycles mentioned previously require more than constant space.
-However, the existence of models such as Population Protocols [aspnes2009population] and Stone Age Computing [emek2013stoneage] which only use constant space demonstrates that such constraints do not preclude the existence of solutions to non-trivial problems.
-The main problem that we consider in the following sections is Leader Election [LeLann77], which is an important primitive in distributed systems as it allows the coordination of processes in the system.
-The goal in this problem is for a unique node to enter and remain in the elected state indefinitely.
-A usual assumption within this context, that we do not make in this work, is that nodes have distinct identifiers.
+As shown initially by [naor1993lcl], there exist distributed problems that admit local solutions (i.e., solutions that can be verified without additional information) for which a certificate is not needed.
+The specification of the set of allowed labelings (i.e., local solutions) to these problems up to some constant radius $r$ are called Locally Checkable Labelings (LCL).
+An example of a distributed problem with a local solution is vertex coloring in which every node knows whether the solution is locally correct by only comparing its label (i.e., its color) to that of its neighbors at distance one.
+The labels we define in this work are an orientation of the edges incident to each node.
+For each class of graphs we consider, we define a set of local conditions and we show that if all conditions we define for each family of graphs hold at every node, the global configuration contains a unique sink that we define to be a leader.
+Then we use this description of what a solution looks like locally to construct a silent self-stabilizing algorithm for leader election in specific classes of graphs that works under a Gouda fair scheduler [Gouda2001theory] which is the strongest kind of fairness [dubois2011taxonomy].
+Our self-stabilizing algorithm is generic and works for any problem for which the solution can be defined as a local labeling.
+Results in a similar flavor that combine local checks [afek1997local] with a global “reset” [awerbuch1994reset] to some predetermined state in order to achieve self stabilization already exist [awerbuch1994stabilization].
+However, contrary to algorithms such as [awerbuch1994reset, awerbuch1994stabilization, BlinFP14], our algorithm only needs one additional bit of memory to transform a labeling scheme to a self-stabilizing algorithm, due to Gouda fairness.
+Leader Election [LeLann77], is an important primitive in distributed systems as it allows the coordination of processes in the system.
+The goal in this problem is for a unique node to enter and remain in the distinct elected state indefinitely.
+A usual assumption within this context, that we do not make in this work, is that each node has a unique identifier.
 In the anonymous setting, the problem cannot be solved without additional assumptions due to symmetries [angluin1980local].
 One way to overcome this impossibility is to characterize graph classes where the problem can be solved (e.g., as in [yamashita1996computing]) or to employ randomization (e.g., [itai1990symmetry]).
-In this work, we use the former method and provide certificates for leader election for the case of chordal and $K_{4}$-free dismantlable graphs.
-Both chordal graphs and dismantlable graphs can be characterized by an ordering of their vertices.
+In this work, we adopt the former approach and study the problem in specific classes of graphs.
+For each class of graphs we define an LCL problem and we show that any labeling satisfying the conditions is a solution to the leader election problem.
+We give a formal definition of the problem for each class of graphs and each set of conditions in Sections 3.1, 3.2 and 4.
+We focus on $K_{4}$-free dismantlable graphs.
+Dismantlable graphs are exactly the class of cop-win graphs in the classic Cops and Robbers game [aigner1984cops, nowakowski1983vertex, quilliot1978jeux] and include well-structured families such as bridged [FaJa, SoCh] and Helly [BP-absolute, BaPr] graphs, which are known for their convexity and metric properties.
+We briefly consider two additional classes: trees (Section 3.1) and chordal graphs (Section 3.2).
+Trees are particular chordal graphs, which in turn are particular dismantlable graphs.
+Both dismantlable and chordal graphs can be characterized by an ordering of their vertices.
 Chordal graphs admit an ordering called a simplicial elimination scheme, in which each vertex is adjacent to a clique of nodes that appear later in the ordering.
 Similarly, dismantlable graphs admit a dismantling order, where each vertex $v$ has a neighbor that is adjacent to all other neighbors of $v$ appearing later in the ordering.
-Dismantlability, in particular, is a property shared by many important classes in metric graph theory, such as bridged [FaJa, SoCh] and Helly graphs [BP-absolute, BaPr].
-In this work, we design certificates that take advantage of these orderings to define and locally verify an orientation of the edges incident to each node, ensuring that the graph contains a unique sink node, which we define to be the leader.
 1.1 Related Work
-In the LOCAL model introduced by Linial [linial1987local, linial1992journal], it is assumed that nodes exchange messages with neighboring nodes until a task is solved.
-The complexity of an algorithm within this model thus depends on the size of the messages and on the number of synchronous communication rounds.
-An alternative definition of the model that is more appropriate for our setting, is to assume that every node can see the state of each node in its neighborhood up to some distance $d$ and based on this partial knowledge of the graph, choose an output.
-The goal in this case is to minimize $d$ and the size of states for some given task.
-What can be computed within this setting has received significant attention (e.g., [vishkin1986determinisitic, goldberg1987parallel, luby1986simple, alon1986fast]).
-The concept of locally verifiable problems was formalized by Naor and Stockmeyer in their seminal paper [naor1993lcl, naor1995journal] in which they defined the family of Locally Checkable Labeling (LCL) problems.
-A problem is in LCL if given the view, consisting of constant-sized input and output labels, at some constant distance $d$ for each node, there exists an algorithm that can decide whether a computed solution is correct.
-Although LCL problems were continually explored since their introduction, no groundbreaking results emerged until 2016 when Brandt et al.
-[brandt2016lower] led to a series of papers (e.g., [balliu2018classes, balliu2018locally, balliu2020randomness]) that gradually provided a complete picture on the classification of the round complexity of all LCL problems in four broad classes [suomela2020landscape].
+The concept of locality in distributed computing was formalized by Linial [linial1987local, linial1992journal] in the LOCAL model as a way to study how local computations can synthesize a solution to a global problem.
+The focus of this model is on the distance from which each node must gather information in order to compute its part of the global solution.
+Hence the complexity of an algorithm in LOCAL depends on the number of synchronous communication rounds.
+What can be computed within this setting immediately received significant attention (e.g., [vishkin1986determinisitic, goldberg1987parallel, luby1986simple, alon1986fast]) and is still an active area of research.
+A closely related model, CONGEST [peleg2000distributed], is defined in the same way as LOCAL with the restriction that nodes can only exchange bounded information in each round, typically logarithmic to the size of the network.
+In the same spirit, Naor and Stockmeyer defined the class of Locally Checkable Labeling (LCL) problems in their seminal paper [naor1993lcl, naor1995journal].
+An LCL problem is specified by a constant radius $r$ and a finite set of allowed labeled radius $r$ neighborhoods such that, in graphs of bounded degree, a labeling is globally correct if and only if every node’s radius $r$ neighborhood belongs to this set.
+Intuitively the definitions of LOCAL and LCL are connected.
+This connection is made explicit in Chang’s [chang2024lcl] definition of LCL: “A distributed problem on bounded-degree graphs is a locally checkable labeling (LCL) if there is some constant $r$ such that the correctness of a solution can be checked locally in $r$ rounds of communication in the LOCAL model”.
+A systematic study of LCL problems was started in Brandt et al.
+[brandt2016lower] which then led to a series of papers (e.g., [balliu2018classes, balliu2018locally, balliu2020randomness]) that eventually formed a complete picture on the classification of the round complexity of all LCL problems in four broad classes [suomela2020landscape].
 Recently, Bousquet et al.
-[feuilloley2025certification] proposed a different direction of classification based on space instead of round complexity.
+[feuilloley2025certification] proposed a different direction of classification of LCL problems based on space instead of round complexity.
 Several results are given in [feuilloley2025certification], characterizing both “gaps” (i.e., pairs of functions $f(n),g(n)$, with $f(n)<g(n)$ where no property has optimal certificate size in $(f(n),g(n))$ for a given class of graphs) as well as proving the existence of properties with optimal certificate sizes for various classes of graphs including paths, cycles and caterpillars.
@@
 Our approach is complementary to that of [feuilloley2025certification] in the sense that we only consider specific problems but for larger classes of graphs.
-The most similar paper to our work is [chalopin2024stabilising].
-Therein, the authors propose a certificate and a set of conditions with constant memory per node, ensuring a unique sink for a simply connected subset of the regular triangular grid, where simply connected means that the configuration does not contain holes.
-Observe that simply connected parts of triangular grids are $K_{4}$-free dismantlable graphs, hence this work (in particular Section˜4) generalizes the result of [chalopin2024stabilising] and extends this idea to additional classes of graphs.
-Additionally, our definition of view is similar to the concept of binoculars introduced in [chalopin2015binoculars].
-Binoculars provide each node $v$ with all information on the graph induced by the neighborhood of $v$ at distance one.
+The most similar paper to our work is [chalopin2024stabilising], which also locally checks a set of conditions that ensure the existence of a unique sink defined to be a leader in a specific class of graphs.
+The main similarity between our work and [chalopin2024stabilising] is the assumption that nodes are anonymous.
+In [chalopin2024stabilising], the authors propose a constant-memory labeling and a set of sufficient conditions to ensure a unique sink for a simply connected subset of the regular triangular grid $S$, where simply connected means that the configuration does not contain “holes” (i.e., subsets of the triangular grid that do not belong in $S$ and are disconnected to each other).
+As simply connected parts of triangular grids are $K_{4}$-free dismantlable graphs, this work (in particular Section˜4) generalizes the result of [chalopin2024stabilising] and extends this idea to additional classes of graphs.
+The natural generalization of [chalopin2024stabilising] would be considering graphs with a contractible clique-complex, but this class is difficult to handle algorithmically, as recognizing these graphs is undecidable (this follows from a result of Novikov, as explained in [Tancer2016collapsible]).
+Dismantlable graphs form a class of graphs that have contractible clique-complexes, but they are easier to handle as they can be recognized in polynomial time [nowakowski1983vertex, quilliot1978jeux], making them a natural class to consider.
+Chordal graphs form a subclass of dismantlable graphs that is already very large as they contain all split graphs and there are $2^{\Theta(n^{2})}$ split graphs [BRW-split1985, oeis-split].
+Observe that adding a universal vertex (i.e., a vertex neighboring to every other vertex in the graph) to any $n$-node graph results in an $(n+1)$-node dismantlable graph implying that the number of dismantlable graphs with $n+1$ vertices is larger than the number of $n$-node graphs.
+If the initial graph is additionally triangle-free, then the resulting graph is $K_{4}$-free and thus the number of $K_{4}$-free dismantlable graphs of size $n$ is $2^{\Theta(n^{2})}$ [EKR-KnFreeGraphs1976, oeis-triangle-free].
+In fact, the class of $K_{4}$-free bridged graphs (which is a subclass of $K_{4}$-free dismantlable graphs is already very general.
+Indeed, Weetman [Weetman1994] shows that starting from any regular graph $G$ where every cycle has length at least 6, one can construct an infinite graph $H$ where the neighborhood of any vertex induces a graph isomorphic to $G$.
+This graph $H$ is an infinite $K_{4}$-free bridged graph and any ball of $H$ (and more generally any finite convex subgraph of $H$) induces a finite $K_{4}$-bridged graph.
+Recently, [feuilloley2025proving] studied the verification aspect of a unique leader existing in anonymous grids and chordal graphs, giving an $O(\log D)$ distance based certification scheme, where $D$ is the diameter of the network.
+After the first version of the current paper appeared, the idea of [feuilloley2025proving] to verify the existence of a unique leader by a certificate encoding the distance of each node to the leader was extended to more classes of graphs in [ChChKo-meshed].
+Notably, [ChChKo-meshed] verifies the existence of a unique leader in meshed graphs using certificates of constant size, where the certificate of each node is the modulo 3 distance to the leader.
+Meshed graphs is a class of graphs that also contains chordal, bridged, and Helly graphs.
+However, the classes of meshed and ($K_{4}$-free) dismantlable graphs are not comparable.
+In fact, meshed graphs are defined by a distance condition whereas dismantlable graphs do not admit such a metric characterization.
+As a result it is not clear whether a similar distance based labeling scheme can be defined for ($K_{4}$-free) dismantlable graphs.
+In this work, the labels we assign at each node encode an orientation that implies the existence of a unique sink, defined to be a leader.
+Hence, the size of our labels depends on the maximum degree of the graph.
+This means that we only obtain a constant size labeling in graphs of bounded degree.
+This work as well as the work of [feuilloley2025proving] and [ChChKo-meshed] assume that nodes have access to information that fits within the radius two verification rules.
+To the best of our knowledge no other papers consider leader election in LCL or related models in anonymous graphs.
 1.2 Our Contributions
-We define constant-size certification schemes for verifying the existence of a unique sink in chordal graphs and $K_{4}$-free dismantlable graphs, under the assumption that each node can see the subgraph induced by its one-hop neighbors and the certificates of its neighbors.
-This notion of view can be seen as an intermediate assumption, requiring more information than standard radius one visibility but without accessing nodes at distance two.
-For chordal graphs, the certification scheme also ensures a valid acyclic orientation, which is particularly important as acyclicity cannot, in general, be verified with constant-size certificates.
-In the case of dismantlable graphs, we also provide a certification scheme to ensure the existence of a spanning tree, assuming there is a distinct node marked as the root.
-The certification scheme verifying the existence of a unique sink in $K_{4}$-free dismantlable graphs can be used to verify the existence of such a root.
-All of our certificates require only constant size per edge.
-In the last section, we show that given a constant-size certificate, a silent self-stabilizing algorithm can always be constructed with constant memory per process, assuming a Gouda fair scheduler.
-Let $k$ be the number of bits needed to verify a task $T$ using a certificate.
-Our algorithm transforms the certificate to a self stabilizing algorithm using $k+1$ bits.
-This is less than prior general transformations, such as the one in [BlinFP14], which requires $O(k+\log n)$ bits, where $n$ is the number of computational entities in the system, but works even under an unfair scheduler.
-As a result, Algorithm 1 can be used in systems with constant memory where [BlinFP14] cannot be employed.
-Our results demonstrate that self-stabilization with strong global guarantees (i.e., unique sink, acyclic orientation) is possible with limited local memory in natural graph classes.
-That is, we show that combining certificate-based verification and Gouda fairness provides a way for designing fault-tolerant systems even when only constant memory is available.
+As already defined, an LCL is a specification describing a set of allowed radius $r$ neighborhoods, where “allowed” refers to conditions proving that the labels solve a global problem.
+An algorithm is then needed as a procedure that produces those labels.
+In this work:
+1.
+We define a set of labels, which are usually edge orientations, and a set of conditions that define the allowed labelings within the graph induced by the one neighborhood for each node in $K_{4}$-free dismantlable graphs and related classes.
+The problem we solve is ensuring that the orientation of all edges incident to every node satisfies a number of conditions that we separately define for each class of graphs.
+We show that when those conditions globally hold, the network contains a unique sink that we define to be a leader.
+The size of the label at each node is deg bits (where deg is the degree of the node).
+As most of our labeling schemes are an orientation of all edges incident to each node, our approach results in constant size labels for graphs of bounded degree or bounded degeneracy [feuilloley2021degeneracy, feuilloley2023genus].
+Although in the distributed verification and LCL literature it is usually assumed that nodes have unique identifiers, here we only consider anonymous nodes.
+Instead of the usual approach of using structural properties for the construction of the labels, in this work we capture the geometry of graphs in the labels.
+This geometric approach can be useful in practical geometric settings, as for example our result holds for planar triangulations where inner nodes have degree at least six.
+As planar graphs are also 5-degenerate, this is an example of a class of graphs with bounded degeneracy where our approach leads to constant-size labels.
+Despite dismantlable graphs (and chordal graphs in Section 3.2) admitting an ordering of their vertices, the edge orientation satisfying our conditions is much more general as we need to show that every labeling satisfying the conditions results in an orientation with a unique sink.
+In fact, our LCL ensures a unique sink despite the orientation of edges allowing large cycles.
+This highlights the fact that with our orientation one cannot use the well known trick of constructing a rooted spanning tree and electing the root.
+In the particular subcase of chordal graphs, our labeling scheme additionally ensures an acyclic orientation.
+For the case of dismantlable graphs, we also provide a labeling scheme to construct a spanning tree, assuming a distinct node marked as the root is given as input.
+The LCL determining a unique sink in $K_{4}$-free dismantlable graphs can be used to calculate such a root.
+2.
+We show that given an LCL, a silent self-stabilizing algorithm can always be constructed with one additional bit of memory per node, assuming a Gouda fair scheduler.
+Although [goos2016locally] already mentions that locally checkable proofs, which include LCLs, can be interpreted as nondeterministic local algorithms, here we formalize this intuition and extend it to defining a self-stabilizing algorithm.
+The one additional bit needed by Algorithm 1 to transform any $k$-bit labeling scheme to a self-stabilizing algorithm is less than prior general transformations, such as the one in [BlinFP14], which requires $O(k+\log n)$ bits, where $n$ is the number of computational entities in the system, but works even under an unfair scheduler.
+So, Algorithm 1 can be used in systems with constant memory where [BlinFP14] cannot be employed.
+In our case, this means that Algorithm 1 produces an edge orientation with a unique sink, for any initially arbitrarily directed (or undirected) $K_{4}$-free dismantlable or chordal graph.
 2 Model and Preliminaries
@@
 For any vertex $v\in V$, let $N(v)$ denote its neighborhood and let $N[v]=N(v)\cup\{v\}$ denote its closed neighborhood.
+The degree of a vertex $v$ is the size of $N(v)$ and is denoted by $\deg(v)$.
 The 1-ball at $v$, denoted by $B_{1}(v)$, is the graph induced by $N[v]$.
@@
 incoming) neighborhood of $v$ in $G$ and refer to nodes in $N^{+}(v)$ as outneighbors of $v$.
-We define the port labeling function to be $\delta=\{\delta_{v}\}_{v\in V}$, where $\delta_{v}:N(v)\rightarrow\{1,2,\ldots,deg(v)\}$ assigns a unique port number to each edge incident to $v$ for every $v\in V$ (i.e., $\delta_{v}$ is a bijection).
+We define the port labeling function to be $\delta=\{\delta_{v}\}_{v\in V}$, where $\delta_{v}:N(v)\rightarrow\{1,2,\ldots,\deg(v)\}$ assigns a unique port number to each edge incident to $v$ for every $v\in V$ (i.e., $\delta_{v}$ is a bijection).
 We write $K_{4}$ to denote a complete graph with four vertices.
 A graph is $K_{4}$-free if it contains no subgraph that is isomorphic to a $K_{4}$.
-We write directed triangle to refer to a cycle of size three.
+We write directed triangle to refer to a directed cycle of size three.
 Let $G=(V,E,\delta)$ be a graph and $S$ be an arbitrary set of states.
@@
 Spanning Tree
-[…]
-[…]
-  • The set of directed edges $\{\vec{vw}\text{ }|\text{ }v\neq r\}$, where $w=\emph{parent}_{G}(v)$, is a directed spanning tree.
-We slightly adapt the definition of local certification of [feuilloley2025certification] to our setting.
+Let $O=\mathbb{N}$.
+Let $\operatorname{parent}_{G}(v)$ be the unique vertex $w\in N(v)$ such that $\delta_{v}(w)=x_{O}(v)$ if $1\leq x_{O}(v)\leq\deg(v)$ and let $\operatorname{parent}_{G}(v)=\bot$ otherwise (e.g., if $x_{O}(v)=0$).
+The set of correct configurations $P$ is the set of configurations $(G,x_{O})$ such that:
+  • $\forall v\in V$, $0\leq x_{O}(v)\leq\deg(v)$,
+  • ! $\exists!\text{ }r\in V$ such that $\operatorname{parent}_{G}(r)=\bot$ (i.e., such that $x_{O}(v)=0$),
+  • The set of directed edges $\{\vec{vw}\mid v\neq r\text{ and }w=\operatorname{parent}_{G}(v)\}$ is a directed spanning tree.
 We assume that the nodes of $G$ do not have identifiers, that is, $G$ is anonymous.
-Let $L$ be a possibly empty set of input labels.
-If $L\neq\emptyset$, we say that $G$ is labeled and we call the mapping from $V$ to $L$ the input function.
-Let $C$ be a non-empty set of states, called the certificate.
-A certificate assignment on $G$ is a mapping $x_{C}:V\rightarrow C$.
-For a given certificate assignment $x_{C}$ and a node $v\in V$, the view of $v$ consists of the 1-ball $B_{1}(v)$ and the restriction of the port numbering function $\delta$, the input function (if defined) and the certificate assignment $x_{C}$ to $B_{1}(v)$.
+Let $I$ be a possibly empty set of inputs and call the mapping from $V$ to $I$ the input function.
+Let $L$ be a non-empty set of states, called the labels.
+A labeling on $G$ is a mapping $x_{L}:V\rightarrow L$.
+For a given labeling $x_{L}$ and a node $v\in V$, the view of $v$ consists of the 1-ball $B_{1}(v)$ and the restriction of the port numbering function $\delta$, the input function (if defined) and the labeling $x_{L}$ to $B_{1}(v)$.
+[^1] The graph induced by the one neighborhood of a node fits within the usual radius 2 verification rules.
 An example of the view of a node is given in Figure˜1.
-Figure 1: Right: A graph with port labels (red) where each node is assigned a certificate in $\{0,1\}^{*}$ (black). Left: The view of the node marked in gray.
-Let $R$ be a set of configurations of the form $(B,x_{C})$ where $B$ is a ball of radius 1 and $x_{C}$ is a certificate assignment defined on the nodes of $B$.
+Figure 1: Right: A graph with port labels (red) where each node is assigned a label in $\{0,1\}^{*}$ (black). Left: The view of the node marked in gray.
+Let $R$ be a set of configurations of the form $(B,x_{L})$ where $B$ is a ball of radius 1 and $x_{L}$ is a labeling defined on the nodes of $B$.
 We refer to $R$ as a set of local configurations.
-A local configuration is said to be correct with respect to $(C,R)$ if for every $v\in V$ the pair $(B_{1}(v),x_{C}|_{N[v]})$ belongs to $R$, where $x_{C}|_{N[v]}$ denotes the restriction of $x_{C}$ to $N[v]$.
-In other words, the local configuration is correct if the labeling in each ball of radius 1 satisfies the local conditions defining a correct configuration $R$.
-We refer to the pair $(C,R)$ as a certification scheme.
-A certification scheme is used to solve a given task $\mathcal{T}=(O,P)$.
-The key property we prove, is that if a configuration is locally correct with respect to the certification scheme, then it is globally correct (i.e., it is a correct configuration for the task).
-Formally, let $out:C\rightarrow O$.
-We say that there exists a certification scheme $(C,R)$ for some task $(O,P)$ in a family of possibly labelled graphs $\mathcal{F}$ if for every $G\in\mathcal{F}$ the following two properties are satisfied [^1] The symbol $\circ$ denotes the composition operator.:
+A configuration is said to be correct with respect to $(L,R)$ if for every $v\in V$ the pair $(B_{1}(v),x_{L}|_{N[v]})$ belongs to $R$, where $x_{L}|_{N[v]}$ denotes the restriction of $x_{L}$ to $N[v]$.
+In other words, the configuration is correct if the labeling in each ball of radius 1 satisfies the local conditions described by the local configurations in $R$.
+We refer to the pair $(L,R)$ as a labeling scheme.
+A labeling scheme $(L,R)$ is of finite type if for any $1$-ball $B$, there is a finite number of labelings of $B$ in $R$, i.e., the set $|\{x\in C^{V(B)}\mid(B,x)\in R\}|$ is finite.
+For example, if $L$ is finite then any labeling scheme $(L,R)$ is finite.
+Observe that a labeling scheme is of finite type if the number of states a node $v$ can take in a correct labeling of its $1$-ball is bounded by a function of its degree $\deg(v)$.
+Note that in the case of bounded degree graphs, a labeling scheme of finite type corresponds precisely to a classical LCL.
+In the remainder of this paper we only consider labelings of finite type [^2] A similar concept of Locally Finite Labelings (LFLs) was recently introduced in [schmid2026lfl]..
+A labeling scheme is used to characterize the solutions to a given task $\mathcal{T}=(O,P)$.
+The key property we prove, is that if a configuration is (locally) correct with respect to the labeling scheme, then it is (globally) correct with respect to the task specification (i.e., it is a correct configuration for the task).
+Formally, there exists $out:L\rightarrow O$, such that there exists a labeling scheme $(L,R)$ for some task $(O,P)$ in a family of graphs (that might initially have inputs) $\mathcal{F}$ if for every $G\in\mathcal{F}$ the following two properties are satisfied [^3] The symbol $\circ$ denotes the composition operator.:
 Existence
-$\forall G\in\mathcal{F},\text{ }\exists(G,x_{C})\text{ s.t. }\forall v\in V,\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R$
+$\forall G\in\mathcal{F},\text{ }\exists(G,x_{L})\text{ s.t. }\forall v\in V,\text{ }(B_{1}(v),x_{L}|_{N[v]})\in R$
 Soundness
-$\forall(G,x_{c})\text{ s.t. }G\in\mathcal{F},\text{ if }\forall v\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R\text{ then }(G,\emph{out}\circ x_{C})\in P$
-In most of our certification schemes, the certificates correspond to an edge orientation.
-For these certificates, we denote the configuration as $\vec{G}$ and we adopt the following encoding for the certificates: the certificate assigned to each node $v\in V$ is composed of $\deg(v)$ bits where each bit encodes the direction of the corresponding incident edge.
+$\forall(G,x_{L})\text{ s.t. }G\in\mathcal{F},\text{ if }\forall v\text{ }(B_{1}(v),x_{L}|_{N[v]})\in R\text{ then }(G,\emph{out}\circ x_{L})\in P$
+In most of our labeling schemes, the labels correspond to an edge orientation.
+For these labels, we denote the configuration as $\vec{G}$ and we adopt the following encoding for the labels: the label assigned to each node $v\in V$ is composed of $\deg(v)$ bits where each bit encodes the direction of the corresponding incident edge.
 We assume that 0 denotes an incoming edge to $v$ and 1 denotes an outgoing edge from $v$.
-The bit at position $i\in\{0,\ldots,\deg(v)-1\}$ of the certificate corresponds to the direction of the edge at port $i$ of $v$.
+The bit at position $i\in\{0,\ldots,\deg(v)-1\}$ of the label corresponds to the direction of the edge at port $i$ of $v$.
 A visualization of this for the graph in Figure˜1 is given in Figure˜2.
-Figure 2: The $i$-th bit of the certificate (black numbers) of each node $v$ corresponds to the direction of the edge incident to port $i$ (red numbers) of $v$, for $i\in\{0,\ldots,deg(v)-1\}$ where 0 denotes an incoming and 1 an outgoing edge.
-On the left we show the direction of the edges and on the right we additionally show the bit of the certificate of each node corresponding to each edge.
+Figure 2: The $i$-th bit of the label (black numbers) of each node $v$ corresponds to the direction of the edge incident to port $i$ (red numbers) of $v$, for $i\in\{0,\ldots,deg(v)-1\}$ where 0 denotes an incoming and 1 an outgoing edge.
+On the left we show the direction of the edges and on the right we additionally show the bit of the label of each node corresponding to each edge.
 For each edge $uv$ where $u$ (resp. $v$) is connected to $v$ (resp. $u$) through port $i$ (resp. $j$), we say that the edge is oriented if the $i$-th bit of $u$ differs from the $j$-th bit of $v$.
-2.2 Warm Up: A Certificate for Leader Election in Trees
-We give a simple certificate that guarantees a unique sink in the class of trees.
-While not new (e.g., one could easily construct a certificate following the self–stabilizing leader election algorithm of [antonoiu1996trees]), we believe it gives a good intuition behind the certificates and proofs in the following sections.
-The certificate of each node consists of the orientation of edges incident to the node.
+3 Warm Up: Labelings for Trees and Chordal Graphs
+3.1 Unique Sink in Trees
+We give a simple labeling that guarantees a unique sink in the class of trees.
+While not new (e.g., one could easily construct a certificate following the self–stabilizing leader election algorithm of [antonoiu1996trees]), we believe it gives a good intuition behind the labelings and proofs in the following sections.
+The label of each node consists of the orientation of edges incident to the node.
 A correct configuration is defined by two conditions, T1 and T2.
@@
 Theorem 3.1 (Soundness).
-In an unlabeled tree $T$, any configuration $\vec{T}$ satisfying T1 and T2 contains a unique sink.
-Proof 2.2.
+In an anonymous tree $T$ where the nodes of $T$ do not have any input, any configuration $\vec{T}$ satisfying T1 and T2 contains a unique sink.
+Proof 3.2.
 We prove the theorem by induction on the number of vertices of $T$.
@@
 Hence, in all cases $\vec{T}$ has a unique sink, completing the proof.
-3 Chordal Graphs
-We generalize the certification scheme from Section 2.2 to ensure the validity of a certificate assignment, which guarantees acyclicity and a unique sink in chordal graphs.
-Each node’s certificate is the orientation of its incident edges.
-We refer to the nodes of $G$ and the certificate assignment as a configuration $\vec{G}$.
+3.2 Unique Sink and Acyclic Orientation in Chordal Graphs
+We generalize the labeling scheme from Section 3.1 to a set of conditions guaranteeing acyclicity and a unique sink in chordal graphs.
+Each node’s label is the orientation of its incident edges.
+We refer to the nodes of $G$ and the labeling as a configuration $\vec{G}$.
 A graph is chordal if every induced cycle is of length three.
@@
     Vertex $v$ does not belong to any directed triangle.
-Observe that the verification of C3 is possible, since the view by definition contains the orientation of edges between neighboring nodes.
+Observe that checking C3 is possible, since the view by definition contains the orientation of edges between neighboring nodes.
 {observation}
-[Existence] A correct configuration with respect to C1–C3 exists and can be constructed as follows: direct all edges incident to a simplicial vertex $u$ toward $N(u)$, then iteratively restrict the graph to the subgraph of remaining unoriented edges and orient edges away from simplicial vertices.
+[Existence] A correct configuration with respect to C1–C3 exists and can be constructed as follows: pick a simplicial vertex $u$, direct all edges incident to $u$ toward $N(u)$, then iteratively restrict the graph to the subgraph of remaining unoriented edges and orient edges away from simplicial vertices.
 Lemma 3.3.
 In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 is an acyclic orientation of $G$.
-Proof 3.2.
+Proof 3.4.
 Suppose the lemma statement does not hold and $\vec{G}$ contains an oriented cycle.
@@
 In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 contains a unique sink.
-Proof 3.4.
+Proof 3.6.
 We prove the lemma by induction on the number of vertices of $G$.
@@
 Since $N(u)$ is a complete graph, this $t^{\prime}$ is unique.
-Furthermore, as we have already shown that $t^{\prime}$ has no outneighbor outside of $N[u]$, $t^{\prime}$ is the unique sink of $\vec{G}\setminus\{u\}$.
-Since $^{\prime}\in N^{-}(u), tt$ is not a sink in $\vec{G}$.
+Furthermore, as we have already shown that $t^{\prime}$ has no outneighbor outside of $N[u]$, $t^{\prime}$ is the unique sink $t$ of $\vec{G}\setminus\{u\}$.
+Since $^{\prime}=t\in N^{-}(u), tt$ is not a sink in $\vec{G}$.
 Therefore, $u$ is the unique sink of $\vec{G}$.
-Theorem 3.5 directly follows from Section˜3 and Lemmas˜3.1 and 3.3.
+Theorem 3.7 directly follows from Section˜3.2 and Lemmas˜3.3 and 3.5.
 Theorem 3.7.
 In a chordal graph $G$, any configuration $\vec{G}$ satisfying C1–C3 is an acyclic orientation of $G$ containing a unique sink.
+Observe that to construct a spanning tree in a chordal graph satisfying C1–C3, it is enough that each non-root node picks an outgoing edge.
 4 Dismantlable graphs
-This section is divided into two parts: a certification scheme ensuring a unique sink in $K_{4}$-free dismantlable graphs and a certification scheme for a rooted spanning tree, given a unique root, in dismantlable graphs.
-Let $G$ be a graph and $u,v$ be two vertices in $V(G)$.
-We say that $v$ is dominated by $u$ in $G$ if $N[v]\subseteq N[u]$.
-A graph is dismantlable if there exists an ordering $v_{1},\ldots,v_{n}$ of $V(G)$ such that in $G[\{v_{1},\ldots,v_{i}\}]$, $v_{i}$ is dominated by some $v_{j}$ with $j<i$.
-So for each $i>1$, there exists $j<i$ such that $N[v_{i}]\cap\{v_{1},\ldots,v_{i}\}\subseteq N[v_{j}]$.
-Dismantlable graphs are exactly the class of cop-win graphs in the classic Cops and Robbers game [aigner1984cops, nowakowski1983vertex, quilliot1978jeux] and include well-structured families such as Bridged graphs and Helly graphs, which are known for their convexity and metric properties.
+[…]
 4.1 Leader election in $K_{4}$-free dismantlable graphs
-In this section we only consider $K_{4}$-free dismantlable graphs.
-As in Section 3, the certificate at each node consists of an orientation on the edges incident to the node and we write $\vec{G}$ to denote a configuration.
-We define a locally correct configuration to be one that satisfies the following rules for every node $v$ of $G$.
-We show that if all rules are satisfied, then $G$ contains a unique sink that we define to be the leader.
+Here we only consider $K_{4}$-free dismantlable graphs.
+As in Section 3.2, the label at each node consists of an orientation on the edges incident to the node and we write $\vec{G}$ to denote a configuration.
+We define a locally correct configuration to be one that satisfies D1–D3 for every node $v$ of $G$.
+We show that if all rules are satisfied, $G$ contains a unique sink.
   1. D1
@@
 Orient each edge between two nodes $v_{i}$ and $v_{j}$ as $\overrightarrow{v_{i}v_{j}}$ if and only if $i>j$ in the dismantling order.
+An example of this orientation for the dismantiling order $(i,h,g,f,e,d,c,b,a)$ of the graph in Figure 3 is given in Figure 3.
 This orientation satisfies D1 and D3.
-For D2, consider any node $v$.
-If $N^{+}(v)=\emptyset$, D2 is immediately satisfied for $v$.
-So suppose $N^{+}(v)\neq\emptyset$.
-By dismantlability, every node $v$ is dominated by some later node $u$ in the dismantling order.
-Then $\overrightarrow{vu}$ is an edge of the orientation and for every other outneighbor $w$ of $v$, $w\in N(u)$.
-Hence, $N^{+}(v)\subseteq N[u]$, satisfying D2.
-However, notice that in the case of $K_{4}$-free dismantlable graphs, one cannot hope to have a property similar to Lemma˜3.1.
-Indeed, as in [chalopin2024stabilising], one can construct configurations satisfying D1-D3 that contain directed cycles.
-An example is shown in Figure˜3.
-Figure 3: A $K_{4}$-free dismantlable graph and a certificate assignment that satisfies D1–D3 but is not acyclic.
-The square node dominates all cyclic nodes.
+For D2, consider any node $v_{i}$.
+If $i=1$, then $N^{+}(v)=\emptyset$, and D2 is immediately satisfied for $v$.
+Suppose now that $i>1$.
+Since $v_{1},v_{2},\ldots,v_{n}$ is a dismantling order, there exists $j<i$ such that $N[v_{i}]\cap\{v_{1},\ldots v_{i}\}\subseteq N[v_{j}]$.
+By the definition of the orientation of $G$, $N^{+}(v_{i})=N(v_{i})\cap\{v_{1},\ldots v_{i}\}$ and thus $N^{+}(v_{i})\subseteq N[v_{j}]$, i.e., D2 holds for $v_{i}$ with $v_{i}^{+}=v_{j}$.
+Notice that in the case of $K_{4}$-free dismantlable graphs, one cannot hope to have a property similar to Lemma˜3.3.
+As in [chalopin2024stabilising], one can construct configurations satisfying D1-D3 that contain directed cycles (e.g., Figure˜3) and are thus not obtained from a dismantling order.
+Figure 3: (a) An example of a $K_{4}$-free dismantlable graph.
+(b) A $K_{4}$-free dismantlable graph and an acyclic labeling that satisfies D1–D3.
+(c) A $K_{4}$-free dismantlable graph and a labeling that satisfies D1–D3 but is not acyclic.
 The edges forming the cycle are marked in orange.
@@
 If $w^{+}=v$, i.e., $N^{+}(w)\subseteq N[v]$, then $N^{+}(w)\subseteq N[v]\subseteq N[u]$ and we can thus choose $w^{+}=u$.
-In the remainder of the proof we show that D2 is satisfied in $\vec{G}^{\prime}$ and that $\vec{G}$ contains a unique sink.
+In the remainder of the proof, we distinguish different cases.
+In each case, we first show that D2 is satisfied in $\vec{G}^{\prime}$, then we apply the induction hypothesis on $\vec{G}^{\prime}$ to establish that $\vec{G}^{\prime}$ contains a unique sink, and finally, we deduce that $\vec{G}$ also contains a unique sink.
 We distinguish two cases, depending on whether $uv$ is directed as $\overrightarrow{uv}$ or $\overrightarrow{vu}$.
@@
 Case 2.1.1.
-$N^{+}(u)\cap N^{-}(v)=\emptyset$
+$N^{+}(u)\cap N^{-}(v)=\emptyset$.
 {claimproof}
@@
 In either case, D2 is satisfied for $u$ in $\vec{G}^{\prime}$.
+Observe also that as $N[v]\subseteq N[u]$ and as $N^{+}(u)=N^{+}[v]$, necessarily we have $N^{-}(v)\subseteq N^{-}[u]$.
 By induction hypothesis $\vec{G}^{\prime}$ contains a unique sink $t$.
 Notice that as $N^{+}(u)=N^{+}[v]$, $v$ is a sink of $\vec{G}$ if and only if $u$ is a sink of $\vec{G}^{\prime}$.
-Therefore, if $t\neq u$ in $\vec{G}^{\prime}$, $v$ is not a sink in $\vec{G}$ and $t$ is the unique sink in $\vec{G}$.
-If $t=u$, then $t$ is not a sink in $\vec{G}$ but $v$ is, consequently $v$ is the unique sink of $\vec{G}^{\prime}$.
+Therefore, if $t=u$, then $t$ is not a sink in $\vec{G}$ but $v$ is, and consequently $v$ is the unique sink of $\vec{G}$.
+Suppose now that $t\neq u$.
+Then $v$ is not a sink in $\vec{G}$ as $N^{+}[v]=N^{+}(u)$ is non-empty.
+Moreover, since $N^{-}(v)\subseteq N^{-}[u]$ and since $t\notin N^{-}(u)$, $v\notin N^{+}(t)$ and $t$ is also a sink in $\vec{G}$.
+Thus, $t$ is the unique sink in $\vec{G}$.
 In either case, $\vec{G}$ contains a unique sink concluding the proof of ˜2.1.1.
 Case 2.1.2.
-$N^{+}(u)\cap N^{-}(v)\neq\emptyset$
+$N^{+}(u)\cap N^{-}(v)\neq\emptyset$.
 {claimproof}
@@
 Therefore, D2 is satisfied at $w$ in $\vec{G}^{*}$.
-Hence D2 holds in $\vec{G}^{*}$.
+So D2 holds in $\vec{G}^{*}$.
 Consequently, by ˜2.1.1 applied to $\vec{G}^{*}$, we deduce that there exists a unique sink $t$ in $\vec{G}^{*}$.
@@
 We first show that D2 is satisfied in $\vec{G}^{\prime}$.
-For D2 to be violated, $v$ should be the only choice for $w^{+}$, for some $w\neq u$.
-If $N^{+}(w)=\{v\}$, D2 trivially holds for $w$ in $\vec{G}^{\prime}$.
-If $u\in N^{+}(w)$, D2 holds for $w$ by Lemma˜4.3.
-Suppose now there exists $z\in N^{+}(w)\setminus\{v\}$.
-Then $z\in N(v)$ and from $u$ dominating $v$, $\{w,z\}\in N(u)$.
-Hence, $v,u,w,z$ induce a $K_{4}$, contradiction.
-So we can assume that $w^{+}\neq v$ for all $w\in V$, and D2 holds for all nodes in $\vec{G}^{\prime}$.
-By induction hypothesis, $\vec{G}^{\prime}$ contains a unique sink $t$. We consider two cases, based on the orientation of $u^{+}v$.
+More precisely, we show that for each $w\neq v$, either $N^{+}(w)\setminus\{v\}=\emptyset$, or there exists $w^{+}\neq v$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $\vec{G}$ (and thus in $\vec{G^{\prime}}$).
+Note that if $N^{+}(w)=\emptyset$, the property holds.
+Otherwise, there exists $w^{+}$ such that $N^{+}(w)\subseteq N[w^{+}]$ in $G$.
+If $w^{+}\neq v$ or if $N^{+}(w)=\{v\}$, we are done.
+Suppose now that $w^{+}=v$.
+Since $w\in N(v)$, we have $w\in N(u)$, and if $u\in N^{+}(w)$, by Lemma˜4.3, we can assume that $w^{+}=u$ and we are also done.
+Suppose now that $w^{+}=v$, that $u\in N^{-}(w)$, and that there exists $z\in N^{+}(w)\setminus\{v\}$.
+Then $z\in N(v)\subseteq N[u]$ and $z\neq u$ as $w\in N^{-}(z)\cap N^{+}(u)$.
+But then $v,u,w,z$ induce a $K_{4}$, which is a contradiction.
+This shows that D2 holds for all nodes in $\vec{G}^{\prime}$.
+By induction hypothesis, $\vec{G}^{\prime}$ contains a unique sink $t$.
+For any vertex $w$ of $\vec{G^{\prime}}$ distinct from $t$, we have just shown that we can assume that $w^{+}\neq v$.
+We consider two cases, based on whether $t=u^{+}$ or not.
+Note that in both cases, $v\in N^{+}(u)\subseteq N[u^{+}]$ and thus $v$ is a neighbor of $u^{+}$.
 Case 2.2.1.
-The edge $u^{+}v$ is $\overrightarrow{vu^{+}}$, that is, $u^{+}\in N^{+}(v)$.
+$t=u^{+}$
 {claimproof}
-If $t$ is not a sink in $\vec{G}$, $t\in N^{-}(v)$ and thus, $t\in N(u)$.
-Observe that it cannot be that $t=u^{+}$ since we have assumed $\overrightarrow{vu^{+}}$, which means that $vt$ is oriented as $\overrightarrow{vt}$ and $t$ would still be a sink in $\vec{G}$ contradicting our assumption.
-So suppose $t\neq u^{+}$.
-Since $t$ is a sink in $\vec{G}^{\prime}$, $t\in N^{+}(u)$ and thus $t\in N(u^{+})$, forming a $K_{4}$ among $u,v,t,u^{+}$.
-Therefore, $t\notin N^{-}(v)$ and $t$ is also a sink in $\vec{G}$.
-Since $v$ is not a sink in $\vec{G}$ due to $\overrightarrow{vu^{+}}$, $t$ is the unique sink in $\vec{G}$.
+Since $u^{+}$ is the unique sink of $\vec{G}^{\prime}$, if $v\in N^{-}(u^{+})$, then $u^{+}$ is the unique sink of $\vec{G}$.
+Suppose now that $v\in N^{+}(u^{+})$ and observe that $v$ is the only potential sink in $\vec{G}$.
+We claim that $v$ is indeed the unique sink in $\vec{G}$.
+Suppose this is not true and that there exists $z\in N^{+}(v)$.
+Note that $z\notin\{u,u^{+}\}$ as $v\in N^{-}(z)\cap N^{+}(u)\cap N^{+}(u^{+})$.
+Since $N(v)\subseteq N[u]$, $z$ is also a neighbor of $u$.
+From D3 applied to the triangle uvzuvz, $z\in N^{+}(u)$ and thus $z\in N[u^{+}]$.
+Consequently, $v,u,u^{+}$ and $z$ induce a $K_{4}$ which is impossible.
 This ends the proof of ˜2.2.1.
 Case 2.2.2.
-The edge $u^{+}v$ is $\overrightarrow{u^{+}v}$, that is, $u^{+}\in N^{-}(v)$.
+$t\neq u^{+}$.
 {claimproof}
-We claim that $u^{+}=t$.
-Suppose for contradiction that $u^{+}\neq t$.
-Then, since D2 is satisfied in $\vec{G}$, there exists $u^{++}\in N^{+}(u^{+})$ such that $N^{+}(u^{+})\subseteq N[u^{++}]$.
-Since we have assumed that $v\in N^{+}(u^{+})$, it must be that $v\in N[u^{++}]$.
-Suppose $v\neq u^{++}$, that is, $v\in N(u^{++})$.
-Then since $v$ is dominated by $u$, the node $u^{++}\in N(u)$.
-However, this means that the vertices $v,u,u^{+}$ and $u^{++}$ induce a $K_{4}$, hence this case is not possible.
-So suppose $v=u^{++}$.
-Suppose that $u^{+}$ has some outneighbor $w$ distinct from $v$.
-Then, since $v=u^{++}$, $w$ must be in $N(v)$ and $u$ is connected to $w$ as $u$ dominates $v$.
-In this case, the vertices $v,u,u^{+}$ and $w$ induce a $K_{4}$ so this setting is not possible and it must be that $N^{+}(u^{+})=\{u^{++}\}=\{v\}$.
-Therefore, it must be that $u^{+}=t$ in $\vec{G}^{\prime}$, contradicting the initial assumption.
-Consider now $u^{+}=t$ in $\vec{G}^{\prime}$, implying that $v$ is the only outneighbor of $u^{+}$ in $\vec{G}$.
-Let $z\in N(v)$.
-Then, $z$ must also be neighboring to $u$ as $u$ dominates $v$.
-We claim that $v$ is the unique sink in $\vec{G}$.
-The orientation $\overrightarrow{uz}$ means that the edge $u^{+}z$ must also exist implying that $u,v,u^{+}$ and $z$ induce a $K_{4}$ which is not possible.
-So $uz$ must be oriented as $\overrightarrow{zu}$.
-From D3, $zv$ should be oriented as $\overrightarrow{zv}$.
-Hence every neighbor of $v$ in $\vec{G}$ must be incoming to $v$.
-Therefore, $v$ is a sink in $\vec{G}$ and since the unique sink $u^{+}$ of $\vec{G}^{\prime}$ is not a sink in $\vec{G}$ due to $\overrightarrow{u^{+}v}$, $v$ is the unique sink in $\vec{G}$.
+If $t$ is not a sink in $\vec{G}$, $t\in N^{-}(v)$ and thus, $t\in N[u]$.
+Since $t$ is a sink in $\vec{G}^{\prime}$, $t\neq u$ as $u$ has an outneighbor $u^{+}$ in $\vec{G}^{\prime}$ and $t\in N^{+}(u)$.
+Since $v,t\in N^{+}(u)\subseteq N[u^{+}]$ and since $u,v,u^{+},t$ are four different vertices, they induce a $K_{4}$ in $G$, which is a contradiction.
+Consequently, $t$ is a sink in $\vec{G}$.
+Suppose now that $t$ is not the unique sink of $\vec{G}$, i.e., that $v$ is also a sink of $\vec{G}$.
+This implies that $u,u^{+}\in N^{-}(v)$.
+Since $u^{+}$ is not a sink in $\vec{G^{\prime}}$, it is not a sink in $\vec{G}$ either and thus there exists $u^{++}\in V(G)$ such that $N^{+}(u^{+})\subseteq N[u^{++}]$.
+Since we can assume that $u^{++}\neq v$ and since $v\in N^{+}(u^{+})\subseteq N[u^{++}]$, we have $u^{++}\in N(v)\subseteq N[u]$.
+Note that $u^{++}\neq u$ since $u^{+}\in N^{+}(u)\cap N^{-}(u^{++})$ Therefore $v,u,u^{+},u^{++}$ induce a $K_{4}$, which is a contradiction.
+This establishes that $t$ is the unique sink of $\vec{G}$.
 This ends the proof of ˜2.2.2.
@@
 4.2 Spanning Tree
-We present a certificate verifying the existence of a spanning tree in any dismantlable labeled graph $G=(V,E)$.
-The labeling function assigns each $v\in V$ a label in $L=\{0,1\}$, with exactly one node labeled 1 (the root $r$) and all others labeled 0.
-We define $C=\{0,1,2\}$ to be the certificate and each node is assigned a value from $C$ by $x_{C}:V\rightarrow C$.
-The following must hold for every node $v$ of $G$ with value $i$.
+We present a labeling constructing a spanning tree in any dismantlable graph $G=(V,E)$, with inputs.
+The input function assigns each $v\in V$ an input in $I=\{0,1\}$, with exactly one node receiving 1 (the root $r$) and all others receiving 0.
+We define $L=\{0,1,2\}$ to be the set of labels and each node is assigned a value from $L$ by $x_{L}:V\rightarrow L$.
+The following must hold for every node $v$ of $G$ with label $l$.
   1. N1
-    Node $v\neq r$ has at least one neighbor with value $(i-1)\mod 3$.
+    Node $v\neq r$ has at least one neighbor with label $(l-1)\mod 3$.
   2. N2
     Node $v$ is not in any three-node triangle labeled $\{0,1,2\}$.
-We show that if N1–N2 are satisfied for every node in $G$, adding an edge from any node $v$ distinct from the root with value $i$ to any neighbor of $v$ with value $i-1\mod 3$, induces a spanning tree.
+We show that if N1–N2 are satisfied for every node in $G$, adding an edge from any node $v$ distinct from the root with label $l$ to any neighbor of $v$ with label $l-1\mod 3$, induces a spanning tree.
+We first prove the Existence property.
 Lemma 4.5 (Existence).
-In any graph $G$ there exists a certificate assignment that satisfies N1–N2 for every node in $G$.
+In any graph $G$ there exists a labeling that satisfies N1–N2 for every node in $G$.
 Proof 4.6.
@@
 In either case, the existence of $\mathsf{C}^{\prime}$ contradicts $\mathsf{C}$ being a cycle with minimum sum of orders, proving the lemma.
-Let $T$ be a spanning subgraph of $\vec{G}$ in which each node $v$, except the root, arbitrarily selects a parent among its neighbors with value equal to $(\textit{value of }v-1)\mod 3$.
+Let $T$ be a spanning subgraph of $\vec{G}$ in which each node $v$, except the root, arbitrarily selects a parent among its neighbors with label equal to $(\textit{label of }v-1)\mod 3$.
 We show that $T$ is a tree.
-Since $\vec{G}$ is acyclic by Lemma 4.7, it follows that $T$ is also acyclic.
-We prove that $T$ is connected.
+Since $\vec{G}$ is acyclic by Lemma 4.7, $T$ is also acyclic.
+We prove $T$ is connected.
 Lemma 4.9.
@@
 Since from Lemma 4.7, $\vec{G}$ and as a result $T$ does not contain cycles, the directed path starting at $u$ must end at a node without a successor in $T$.
-From N1, the root is the only node without a smaller neighbor.
+From N1, the root is the only node without a successor.
 Hence, any node $u$ must be in a path that reaches the root, and as a result $T$ is connected.
-Finally, since every node in $G$ has a label and must be in $T$, from Lemmas 4.7 and 4.9 we obtain the following:
+Since every node in $G$ has a label and must be in $T$, from Lemmas 4.7 and 4.9 we obtain:
 Theorem 4.11 (Soundness).
 In a dismantlable graph, if N1–N2 are satisfied for every node, $T$ is a spanning tree.
-5 From Certificate to Self–Stabilization
-An algorithm is called self-stabilizing when it can autonomously recover from transient faults.
-In particular, a distributed self-stabilizing algorithm is usually defined as a set of rules evaluated at every node.
-Each rule is of the form:
+5 From Labeling to Self–Stabilization
+A distributed self-stabilizing algorithm can autonomously recover from transient fault and is usually defined as a set of rules evaluated at every node. Each rule is of the form:
 $\texttt{name-of-rule}:\langle guard\rangle\longrightarrow command$
@@
 In this setting, Gouda fairness ensures that any configuration that is infinitely often reachable is eventually reached.
-In this section we show how to obtain a silent self–stabilizing algorithm that works under a Gouda fair scheduler, given a certification scheme.
-We say that under a given algorithm $A$, a certificate assignment is final if no node satisfies any of the guards in $A$.
-We define a correct final certificate assignment to be one that is both final with respect to an algorithm $A$ and correct with respect to a certification scheme.
-Call the set of states in a given certificate $C$.
-Informally, every node that detects that a condition of the certification scheme is violated sets its state to a reset state that we denote $\bot$.
-Additionally, every node that has at least one neighbor that has detected an error and reset its state, also changes its state to the reset state.
+In this section we show how to obtain a silent self–stabilizing algorithm that works under a Gouda fair scheduler, given a labeling scheme.
+We say that under a given algorithm $A$, a label assignment is final if no node satisfies any of the guards in $A$.
+We define a correct final label assignment to be one that is both final with respect to an algorithm $A$ and correct with respect to a labeling scheme.
+Call $L$ the set of states in a given labeling.
+Informally, every node that detects a condition of the labeling scheme being violated sets its state to a reset state that we denote $\bot$.
+Every node that has at least one neighbor in a reset state, also changes its state to the reset state.
 Hence, an error detection is propagated throughout the graph from node to neighboring node.
-When a node that is at the reset state is activated, it non-deterministically switches to a state from $C$.
+When a node that is at the reset state is activated, it non-deterministically switches to a state from $L$.
 Gouda fairness ensures that the graph eventually reaches a correct global configuration.
-A node executing Algorithm 1 evaluates the guards in order. If the first guard is enabled, the node performs the corresponding command and deactivates without evaluating the second guard.
-Input: The certification scheme $(C,R)$ and a certificate assignment $x_{C}:V\rightarrow C$
+A node executing Algorithm 1 evaluates the guards in order. If the first guard is enabled, the corresponding command is performed and the second guard is not evaluated.
+Given a labeling scheme $(L,R)$, a graph $G$ and a vertex $v\in G$, a possible state for $v$ in $L$ is a state $s$ from $L$ such that there exists a 1-ball $(B,x)\in R$ where the center of $B$ is labeled by $x(v)=s$ and $B$ is isomorphic to $B_{1}(v)$.
+If $(L,R)$ is a certification scheme of finite type, then the set of possible states for each node in any graph $G$ is finite.
+Input: The labeling scheme $(L,R)$ and a label assignment $x_{L}:V\rightarrow L$
 1exEvery time node $v$ is activated:
-/* states from $C$ are always chosen non-deterministically */
+/* states from $L$ are always chosen non-deterministically */
 S1 : v.state $\neq\bot$ $\land$ $((B_{1}(v),x_{C}|_{N[v]})\notin R$ $\lor$ $\exists u\in N(v):$ u.state = $\bot)$ $\rightarrow$ v.state $\leftarrow\bot$
-S2 : v.state = $\bot$ $\rightarrow$ v.state $\leftarrow$ state from $C$
-Algorithm 1 General Certificate to Self-Stabilizing Algorithm
+S2 : v.state = $\bot$ $\rightarrow$ v.state $\leftarrow$ possible state from $L$
+Algorithm 1 General Labeling to Self-Stabilizing Algorithm
 Theorem 5.1.
-Let $(C,R)$ be a certification scheme satisfying the Existence and Soundness conditins for some task $\mathcal{T}$ in a connected graph $G$.
+Let $(L,R)$ be a finite labeling scheme of finite type satisfying the Existence and Soundness conditions for some task $\mathcal{T}$ in a connected graph $G$.
 Starting from any arbitrary configuration, any Gouda fair execution of Algorithm 1 on $G$ stabilizes to a correct final configuration.
@@
 Let $G^{*}$ be a configuration where no node satisfies the guard of S1 or of S2, that is, $G^{*}$ is a final correct configuration.
-By definition, $\mathcal{T}$ is solved in $G^{*}$, since $(C,R)$ is a certification scheme for $\mathcal{T}$.
-Suppose for contradiction that there exists a configuration $G^{\prime}\neq G^{*}$ that is infinitely often reachable during the execution of Algorithm 1.
-Since $G^{*}$ is final, it cannot be reachable from $G^{\prime}$.
-As $G^{\prime}$ is not final, either $G^{\prime}=G_{\bot}$ (i.e., all nodes are in state $\bot$) or the guard of S1 is enabled for some node.
-In the former case, $G_{\bot}$ trivially appears infinitely often.
-In the latter, by S1, there exists an activation sequence that leads to $G_{\bot}$.
-By Gouda fairness, for every a configuration that appears infinitely often, every reachable configuration also appears infinitely often.
-Hence, in either case $G_{\bot}$ appears infinitely often.
-Due to S2, once all nodes in $G_{\bot}$ are activated, the configuration becomes $G^{*}$, contradiction.
-Thus, no configuration $G^{\prime}\neq G^{*}$ can appear infinitely often.
-A self–stabilizing algorithm can be obtained by combining the respective certification scheme of each of the previous sections with Algorithm 1.
+By the soundness property of $(L,R)$ for $\mathcal{T}$, $\mathcal{T}$ is solved in $G^{*}$.
+Consequently, in any final configuration, $\mathcal{T}$ is solved.
+We now show that any execution eventually reaches a final configuration.
+Since $(L,R)$ is of finite type, for each node $v$, the set of possible states is finite and thus in any execution, every node $v$ enters only in a finite number of different states.
+Since $G$ is finite, the number of different configurations existing in any execution (finite or infinite) is also finite.
+By the existence property of $(L,R)$ for $\mathcal{T}$ in $G$, there exists a configuration $G^{*}=(G,x_{L})$ of $G$ such that in $G$, every 1-ball $(B_{1}(v),x_{L}|_{N[v]})$ belongs to $R$.
+Suppose there exists an infinite execution of the algorithm.
+This execution does not reach any final configuration, i.e., at each step, there is a node that can execute S1 or S2.
+In the second case, it means that there is node that is labeled by $\bot$.
+In the first case, it means that one can reach in one step a configuration containing at least one node labeled $\bot$.
+Starting from a configuration containing one node labeled $\bot$, there exists a sequence of activation of nodes that enables reaching the configuration $G_{\bot}$ where all nodes are in state $\bot$.
+Then by Gouda fairness, in any infinite execution, the configuration $G_{\bot}$ is reached infinitely often.
+From $G_{\bot}$, there exists a sequence of node activations that enables reaching the configuration $G^{*}$.
+Again, by Gouda fairness, this implies that any infinite execution eventually reaches $G^{*}$.
+Since $G^{*}$ is a final configuration where no node can be activated, we obtain a contradiction with the existence of an infinite execution.
+A self–stabilizing algorithm can be obtained by combining the respective labeling scheme of each of the previous sections with Algorithm 1.
 6 Discussion and Open Problems
-In Sections˜3 and 4 we gave certification schemes for leader election, with constant size per edge.
-A natural direction for further research is studying whether our certification scheme from Section˜4 also works for dismantlable graphs.
+In Section˜4 we gave a labeling scheme that implies a unique sink defined to be a leader, with labels encoding the orientation of all edges incident to every node.
+Observe that with our labeling schemes for leader election in chordal graphs and in trees, any node in the graph can be the sink and thus designated as a leader.
+Indeed, any chordal graph $G$ contains at least two simplicial vertices, and thus for any vertex $v$, there exists a simplicial ordering that ends at $v$.
+By Section˜3.2, there exists then an orientation of $\vec{G}$ satisfying C1–C3 where $v$ is the unique sink of $\vec{G}$.
+This property is however no longer true for dismantlable graphs.
+Consider for example the graph $G$ of Figure 3 and observe that $a$ is the only dominated vertex of $G$.
+This implies that $a$ cannot be the sink of any orientation satisfying D1–D3 obtained by Section˜4.1.
+In fact, one can show that in any orientation $\vec{G}$ of $G$ satisfying D1–D3, $a$ is not a sink.
+Indeed, suppose there exists such an orientation.
+From D2 at $b$ and $c$, the edges $bd$ and $ce$ should be oriented as $\overrightarrow{db}$ and $\overrightarrow{ec}$.
+Consider now the edge $de$ and assume without loss of generality that it is oriented as $\overrightarrow{de}$.
+Then $d$ has two non-adjacent outgoing neighbors $b$ and $e$ that do not have a common neighbor, and thus D2 does not hold at $d$.
+We believe however that our labeling scheme can be easily modified to elect any node in the graph, by adding an extra component to the labels of the nodes in order to encode a directed path from the unique sink to the designated leader.
+A natural direction for further research is studying whether our labeling scheme from Section˜4 ensures the existence of a unique sink in any dismantlable graph.
 In particular, only the proof of ˜2.2 depends on the graph being $K_{4}$-free.
 However, removing this restriction has been very challenging.
-If this is not possible, another direction is considering bridged or Helly graphs, which are dismantlable but enjoy interesting properties that could be useful for addressing our problem.
-Alternatively, if our certification scheme can be proven to also work on dismantlable graphs, it is worth exploring other relevant classes.
-For example, [fieux2020hierarchy] gives a hierarchy on classes (based also on the work of [mazurkiewicz2007locally]) that have dismantling orderings, called s-dismantlable, where 1-dismantlable graphs correspond to the dismantlable graphs defined in Section 4.
+One can also try to adapt to ($K_{4}$-free) dismantlable graphs the approach used in [ChChKo-meshed] to design labeling schemes for leader election in meshed graphs.
+However, this also seems very challenging, as it relies heavily on properties of the distance function in meshed graphs, and dismantlable graphs do not enjoy similar properties (even in the $K_{4}$-free case).
+If one can find a labeling scheme for leader election in dismantlable graphs, it is worth exploring other relevant classes.
+For example, [fieux2020hierarchy] gives a hierarchy on classes (based also on the work of [mazurkiewicz2007locally]) that have dismantling orderings, called s-dismantlable, where 1-dismantlable graphs correspond to the dismantlable graphs.
 Informally, for $s>1$, [fieux2020hierarchy] defines a node to be $s+1$ dismantlable if the graph induced by its open neighborhood is $s$-dismantlable and a graph to be $s$-dismantlable if it is reducible to a single vertex by successive deletions of $s$-dismantlable vertices.
-These directions for future research are summarized in Figure˜4.
+For a summary, see Figure˜4.
 Figure 4: A graph class hierarchy, where each arrow represents an inclusion relation.
-The classes highlighted in green are an easy extension of [chalopin2024stabilising] and those in yellow are considered in this paper for leader election certification with constant memory per edge.
-Another potential direction for future research is exploring whether it is possible to automatically transform a certificate into a self-stabilizing algorithm with a weaker scheduler, as defined in the taxonomy of [dubois2011taxonomy], even using non-constant memory.
+The green classes are the result of [chalopin2024stabilising] and subclasses of the graphs in this paper that generalize the result of [chalopin2024stabilising].
+Yellow classes are considered in this paper.
+Blue classes are considered in [ChChKo-meshed].
+Finally, a potential direction for future research is exploring whether it is possible to automatically transform a labeling into a self-stabilizing algorithm with a weaker scheduler, as defined in the taxonomy of [dubois2011taxonomy], even using non-constant memory.
 For example, the transformation from [BlinFP14], which uses logarithmic memory but works even with an unfair scheduler, contrasts with our algorithm, which uses constant memory but requires a Gouda fair scheduler.